### v1

Abstract
Vertical Federated Learning (VFL) is an emergent distributed machine learning paradigm wherein owners of disjoint features of a common set of entities collaborate to learn a global model without sharing data. In VFL, a host client owns data labels for each entity and learns a final representation based on intermediate local representations from all guest clients. Therefore, the host is a single point of failure and label feedback can be used by malicious guest clients to infer private features. Requiring all participants to remain active and trustworthy throughout the entire training process is generally impractical and altogether infeasible outside of controlled environments. We propose Decoupled VFL (DVFL), a blockwise learning approach to VFL. By training each model on its own objective, DVFL allows for decentralized aggregation and isolation between feature learning and label supervision. With these properties, DVFL is fault tolerant and secure. We implement DVFL to train split neural networks and show that model performance is comparable to VFL on a variety of classification datasets.
1 Introduction
Federated Learning McMahan et al. (2017), or FL, was introduced by Google researchers as a strategy for distributed learning, addressing communication efficiency and data privacy. Distributed participants in FL training do not expose their data to any other party. Instead, they train local models on their data, guided by a coordinator that has no knowledge of the agents’ data. Thus, FL extends collaborative machine learning to contexts where data communication is undesirable or forbidden. Recent advances in communication systems and cloud computing infrastructure combined with an increased public awareness and legislation regarding data privacy (such as the EU’s GDPR) have contributed to FL’s advancement as an emergent machine learning paradigm.
Among FL’s developments is Vertical Federated Learning (VFL). Standard FL can be said to have a “horizontal” or “sample-parallel” division of data, i.e. each participant holds unique samples within a shared feature space. On the other hand, VFL participants hold unique features of a common sample space, which is a “vertical” division, and is “feature-parallel”. This allows guest agents who individually have incomplete information about their target to learn meaningful joint representations without sharing data. A typical example: an imaging center, pathology lab and general hospital may come together to predict rare types of cancer, even though health regulations such as HIPAA may disallow them from collecting each others’ data.
VFL guest agents undertake an “entity alignment” phase before training, where they ensure their datasets’ horizontal indexing is identical, i.e., all participants process features belonging to the same entity in any round of training. At each round, each guest passes its features to its local “guest model.” Guests share their local representations with a host. For example, a guest training a vertically federated ANN may send its activations. A boosted decision tree guest may send a weak learner. The host owns data labels and is aware of the identity of the sample or batch that the guests are processing at any time. The host also possesses a “host model” capable of aggregating these representations and making predictions against the labels. The host calculates a loss based on its aggregation and provides feedback (e.g. gradients) to the guests, which then begin the next round of training. We refer to the combined system of guest and host models as VFL models.
VFL has found increasing success in recent years, with various real-world implementations and proposals Liu et al. (2023). However, VFL suffers from some flaws which make its implementation impractical where participants are few in number and each computationally powerful (as in cross-silo scenarios), and entirely infeasible otherwise. One such flaw is that the link between the guests and the host is a single point of failure. Training a batch of data is contingent on the exchange of intermediate results between the guests and the host. If the host waits indefinitely for all guests to communicate, a guest crashing or losing connection would completely halt training. If the host sets a deadline, some inputs to its model will remain unfulfilled. (In the latter case, a late message will also cause a catastrophic fault). If gradients from the host to the guests are lost or arrive in the incorrect order, guest model convergence is hampered. One might dismiss such events as infrequent or at least controllable in cross-silo settings. However, that is naive. When some of the guests (and even the host) could be geographically sparse and/or low-power devices, it is catastrophic.
A fundamental limitation of VFL is imposed by the size of the joint sample space. For instance, databases in groups of large nationwide financial institutions may share many but not all customers. Guests must discard data outside the intersection, which could have been used to learn better guest models. In systems where the intersection is small, there may not be enough samples to train a VFL model altogether. In general, this also hinders VFL’s applicability to systems with large-scale participation. Consider again the hospital example. Suppose the task is now to predict bone marrow cancer. Along with the pathology lab, the imaging center and the general hospital, we also include data collected by an orthopaedic hospital. Even if the overlap of patients at the other facilities is significant, their overlap with patients of the orthopaedic hospital may be too small.
Another drawback of VFL (and FL in general) is the risk of inference attacks from curious guests. Intuitively, a host makes predictions based on information it receives from all guests. The updates or feedback it returns is derived from this prediction and the ground truth. Therefore, feedback sent to one guest contains some mutual information with respect to the others and to the label. Luo et al. (2021); Fu et al. (2022) show that a malicious agent can engineer its messages to the host such that the label and the other guests’ features can be conveniently extracted from the gradients it receives in return. Attempts to safeguard VFL against inference attacks are generally based on differential privacy, homomorphic encryption, and multi-party techniques such as audits and shared secrets Yang et al. (2023); Khan et al. (2023). These operations incur a high computational cost, which might not be feasible outside some cross-silo scenarios. Further, these methods hamper model convergence.
Building intricate mechanisms to mitigate these issues inflates the complexity of the problem. In this work, we attack the source directly. Backpropagating gradients to guests allows inference attacks. The wait for a response or feedback causes synchronization issues. What if we eschewed end-to-end backpropagation (BP) altogether? We present Decoupled Vertical Federated Learning (DVFL), a novel strategy for ANNs to address these shortcomings. The underlying principle is blockwise learning, where gradient-isolated groups of layers of an ANN are trained on local objectives. Such greedily trained ANNs have been shown to perform comparably to those trained by BP Belilovsky et al. (2019); Journé et al. (2023). We decouple guest models from host models by training them on their own local objectives. Similarly, we decouple supervision from aggregation by training host models on local objectives. We use a final transfer learning model to make predictions against labels. This yields two useful system properties. One, this allows redundant hosts. Standard VFL would require hosts to share labels with each other, which is forbidden. Two, we may use separate datasets for feature extraction and label inference.
With redundancy, even if a guest’s connection to one host fails, training may continue as normal on others. If hosts may fail, the redundancy also affords some degree of tolerance. By using separate datasets for inference and representation learning, we may train guest and host models on data which is not necessarily labeled and not necessarily a part of a joint sample space. This makes DVFL applicable to systems with small intersections of sample spaces. Further, the label owner’s supervised model can be trained completely offline, without the risk of faults.
In Section 2, we provide a formal description of VFL and some algorithms which implement it for various machine learning models. We also elaborate on the shortcomings we hope to address with DVFL. In Section 3, we review some closely-related lines of research and use them to motivate DVFL. In Section 4, we provide a formal description of DVFL with an implementation for ANNs. We show some baseline results on benchmark machine learning tasks in Section 5. In the Appendix, we present more results and discuss our design choices. The main contributions of this work are:
  1. 1.
    We present DVFL, a strategy which eschews host-to-guest feedback as well as label-to-host feedback and therefore has inbuilt redundancy, security and fault handling (including faults during the backward pass) while performing comparably to VFL.
  2. 2.
    We show host redundancy improves model performance, especially when there are faults.
  3. 3.
    We demonstrate how DVFL guests might exploit (possibly unlabeled) data outside a limited sample intersection to learn strong representations.
2 Background: Standard VFL
Entities $\mathbf{x}$ are realizations of a fixed distribution $X$ (say). The rows of a dataset are features of a unique entity $\mathbf{x}$. We may serialize these entities using a sample index. Let us say the set of sample indices corresponding to entities the host has labels for is $\mathcal{S}_{\mathrm{labels}}$. Let us say the entity with index $j$ is $\mathbf{x}_{j}$.
2.1 The VFL Hierarchy
Guests: $\mathcal{G}$ is the set of feature-owning collaborators: the guests. ||$\cdot$|| ⋅ | denotes cardinality. Each $g_{i}\in\mathcal{G}$ has its own
  1. 1.
    Dataset $\mathcal{D}_{i}=\left\{\mathbf{x}_{j,\mathcal{F}_{i}}\in\mathbb{R}^{|\mathcal{F}_{i}|}:j\in\mathcal{S}_{i}\right\}$, where $\mathcal{F}_{i}\subseteq\{1,2,\cdots,D\}$ is the subset of $D$ features that the guest owns, and $\mathcal{S}_{i}$ is the set of entities $g_{i}$’s dataset has records for. $\mathbf{x}_{j,\mathcal{F}_{i}}$ are the features $\mathcal{F}_{i}$ of $\mathbf{x}_{j}$.
  2. 2.
    Guest model $m_{i}(\text{ }\cdot\text{ };\theta_{i}):\mathbb{R}^{|\mathcal{F}_{i}|}\to\mathbb{R}^{k_{i}}$ with parameters $\theta_{i}$.
The guest model $m_{i}$’s task is to concisely represent the features $\mathcal{F}_{i}$ of $\mathbf{x}_{j}$, for all sample indices $j$ that are in the intersection of all guests, i.e. $\forall j\in\mathcal{S}_{\mathrm{guests}}$, where $\mathcal{S}_{\mathrm{guests}}:=\bigcap_{i=1}^{|\mathcal{G}|}\mathcal{S}_{i}$ and pass them to the host $h$. Guests do not have access to other guests’ model parameters or data features. In general, they do not have access to labels, either.
Host: The host agent $h$ executes every round of training. The host has access to target labels. The host, given representations for all available features of $\mathbf{x}_{j}$, may use its own model to make useful predictions for $\mathbf{x}_{j}$. The host has
  1. 1.
    Labels $L=\left\{\mathbf{y}_{j}\in\mathbb{R}^{\text{out}}:j\in\mathcal{S}_{\mathrm{labels}}\right\}$:
    • $\mathcal{S}_{\mathrm{labels}}$ is the set of sample indices for which the host has the corresponding labels
    • We assume that the intersection of all the participants’ indices is not null to allow for training, i.e., $\left|\mathcal{S}_{\mathrm{guests}}\cap\mathcal{S}_{\mathrm{labels}}\right|>0$.
  2. 2.
    Host model $m_{h}(\text{ }\cdot\text{ };\theta_{h}):\mathbb{R}^{k_{1}}\times\mathbb{R}^{k_{2}}\times\dots\mathbb{R}^{k_{|\mathcal{G}|}}\to\mathbb{R}^{\text{out}}$ with parameters $\theta_{h}$.
It is possible that $h\in\mathcal{G}$ i.e. the host may also be one of the guests. In such situations, the host is often known as the “active party” and guests as “passive parties.”
If the host has access to the full computational graph, it may adjust all parameters to better predict label $\mathbf{y}$, usually by minimizing the expected value of some loss function. Otherwise, it may update only its parameters $\theta_{h}$ and send intermediate results (e.g. gradients) via which the guests may update their own models. Together, the guests and host solve the following joint optimization problem:
$\theta^{*}=\operatorname*{arg\,min}_{\theta}\mathbb{E}_{\mathbf{x}_{j}\sim X}\left[\ell(m_{h}\left(\hat{\mathbf{x}}_{j,1},\dots,\hat{\mathbf{x}}_{j,{|\mathcal{G}|}});\theta_{h}\right),\mathbf{y}_{j})\right]\,,$
where $\hat{\mathbf{x}}_{j,i}\equiv m_{i}(\mathbf{x}_{j,\mathcal{F}_{i}};\theta_{i})\in\mathbb{R}^{k_{i}}$.
2.2 Private Set Intersection & Entity Alignment
All guests process entities in the same order, each passing the next entity to its local model. For this to happen, all guests must know the set of common entities $\mathcal{S}_{\mathrm{guests}}\cap\mathcal{S}_{\mathrm{labels}}$. This is usually achieved using entity resolution (if required) and some form of Private Set Intersection (PSI).
PSI is a multiparty computation of the intersection of sets belonging to each party without revealing members outside the intersection, typically by exchanging encrypted versions of the set Freedman et al. (2004). Entity resolution is a multiparty computation to generate a set of unique identifiers from features which are not strictly unique or precise. Entity resolution may be performed by comparing cryptographic long keys Hardy et al. (2017); Nock et al. (2018). After the host obtains the results of PSI, there is a phase of entity alignment. In this phase, the host decides the order of entities which all guests must follow and informs them.
2.3 Split Training for Neural Networks
A faithful implementation of VFL for ANNs is SplitNN Gupta & Raskar (2018). SplitNN was originally proposed as an algorithm to distribute the training of a global neural network across two agents by splitting the network depth-wise into two. At the “split layer”, forward and backward signals are communicated between the two devices. [^1] One might say that the standard SplitNN addresses a special case of VFL where the host has access only to labels and the (single) guest has access only to features. The original paper Gupta & Raskar (2018) also presents a protocol for training when the features are horizontally distributed across $N$ guests. Ceballos et al. (2020); Romanini et al. (2021) extend SplitNN for vertically partitioned data by vertically splitting the shallower layers (see Figure 1).
Figure 1: Diagram illustrating the distributed training of a neural network under SplitNN, VFL with SplitNN, and DVFL. The two input data for VFL and DVFL are partial features of the same entity.
Effectively, the guest models $m_{i}$ are ANNs and host model $m_{h}$ consists of a concatenation layer $m_{h}^{0}$ followed by an ANN $m_{H}^{1}$.
2.4 Impracticalities in VFL Training
Vertical Federated Learning is contingent on implicit assumptions regarding the reliability, trustworthiness and scale of participation.
Subversion of these assumptions can be catastrophic. A lost guest output or gradient is a single point of failure. The intersection $\mathcal{S}_{\mathrm{labels}}\cap\mathcal{S}_{\mathrm{guests}}$ may not be meaningfully large, especially when there are many guests. And, it is possible for a malicious guest to infer private data from feedback.
Thus far, the application of VFL has been limited to a small number of cross-silo settings where the silo owners are powerful enough to assume reliability and to afford the overhead of privacy guarantees. Further, cross-silo participation is small in number so scalability is not a concern. VFL is altogether infeasible for a large class of systems where devices are many or have low compute and communication bandwidths.
2.4.1 Single Points of Failure
In order to make a prediction on a batch of data, the host model $m_{h}:\mathbb{R}^{k_{1}}\times\dots\times\mathbb{R}^{k_{|\mathcal{G}|}}\to\mathbb{R}^{\text{out}}$ requires inputs from all guest models. If an input is unavailable to the host model, no prediction can be made– a catastrophic fault. Similarly, local models’ updates require gradients from the host, lacking which the model convergence slows down and may fail altogether. We consider the following fault model for a round of training:
  • Guest faults: A guest device may be unable to compute its output, update its model, or communicate.
  • Host faults: The host itself may fail, and be unable to receive inputs or calculate gradients.
  • Connection faults: The communication link between a guest and the host may be dropped during the forward or backward pass.
From a VFL host’s perspective, a guest fault and communication fault manifest as a missing input to its model $m_{h}(\text{ }\cdot\text{ };\theta_{h}):\mathbb{R}^{k_{1}}\times\mathbb{R}^{k_{2}}\times\dots\mathbb{R}^{k_{|\mathcal{G}|}}\to\mathbb{R}^{\text{out}}$ during training. The host may employ one of two strategies:
  1. 1.
    Wait: The host may wait until all input arrives.
    • Under a guest fault or communication fault, the host will remain waiting indefinitely, and training will not continue.
    • If the host polls an unresponsive guest, it may receive an input intended for the next iteration, which will lead to incorrect model evaluation.
  2. 2.
    Timeout: The host may set a deadline for each round.
    • Under a guest fault or communication fault, training will fail since the host model is missing input.
    • If a message is successful but late, the model is still missing input.
    • If the model skips that training round and moves onto the next, it may receive the late message from the previous round, which will lead to incorrect model evaluation.
Therefore, a communication or guest fault leads to a catastrophic failure for training. When the host fails, the guest models are unable to update their parameters. Since the cardinality of the joint training dataset is that of the intersection of participants’ sample spaces $|\mathcal{S}_{\mathrm{guests}}\cap\mathcal{S}_{\mathrm{labels}}|$ (which may be small), losing updates from a single batch of data may be detrimental to model performance.
2.4.2 Data Intersection
Since VFL training is contingent on entity alignment, it is not usable in scenarios where the intersection of entities known to each guest is small. This is often the case when the number of guests is large. It could also be the case that, although $\mathcal{S}_{\mathrm{guests}}$ is reasonably large, labels are not available for all its members, i.e., $\mathcal{S}_{\mathrm{labels}}\cap\mathcal{S}_{\mathrm{guests}}\subset\mathcal{S}_{\mathrm{guests}}$.
2.4.3 Inference Attacks
Fu et al. (2022) demonstrate that it is possible for a malicious agent to directly infer the true class from gradients (e.g. by analyzing their sign) for many common loss functions during training. The authors also show that after VFL training is complete, it is possible for a curious guest to train its own prediction model if it can gain access to a very small number of labels. If a malicious guest inflates its learning rate during training, the aggregator gives more importance to results from its model. This way, a prediction model trained only on the features of the malicious guest is almost identical to the aggregator model itself. A malicious guest equipped with such a prediction head can make accurate guesses for labels without accessing other guests’ features.
Luo et al. (2021); Erdoğan et al. (2022) show that a generative network can be trained by a set of guests to model the distribution of other guests’ features. The authors also show that a malicious guest only needs to make a good guess for labels to directly infer features of other guests in vertically federated linear classifiers and decision trees.
Typical defenses include differential privacy, gradient compression, homomorphic encryption and secret shares for gradients and model features. However, these methods usually have a negative effect on model convergence rate and tightness Khan et al. (2023). Moreover, implementing such cryptographic algorithms is computationally expensive, and the cost increases with the number of participants.
3 Related work
3.1 Fault-Tolerant Distributed Optimization
The impact of redundancy for fault tolerance in distributed systems is well studied. Liu et al. (2021) provide a rigorous analysis of redundancy in the case of distributed optimization, showing that systems which are $2f$-redundant are approximately resilient to $f$ Byzantine agents. The authors also present two fault-tolerant algorithms for horizontally distributed convex optimization. Liu (2021); Bouhata et al. (2022) extensively review approaches to fault tolerance in distributed learning. Most methods require synchronization by an aggregator and incur computational overhead. Further, most methods involve manipulating gradients, which forces a choice between convergence rate and tightness.
For VFL: Few methods exist for to address fault tolerance in VFL. Li et al. (2023) introduce FedVS, an algorithm for systems where guest models are polynomial networks Livni et al. (2014a; b). FedVS uses a form of secret sharing to introduce redundancy in the input data, and can tolerate a fixed number of late or failed guests. However, FedVS and many of its predecessors Shi et al. (2022); Chen et al. (2020a); Hu et al. (2019) only address faults in the forward pass.
3.2 Federated Transfer Learning
Federated Transfer Learning (FTL) is a strategy for systems where there is some overlap in both sample space and feature space. The host owns labels and some features (i.e. it is an active participant) for entities in its sample space. Guests have limited labels for their task, but have some small overlap in feature space and sample space with the host (“aligned samples”).
Liu et al. (2020) present Secure Transfer Federated Learning (STFL), a framework for a single host and a single guest. Both parties learn a common latent space using local feature extractors. Encodings of the passive participant’s unaligned samples in this feature space can be used by the passive participant to learn its task model. Feng & Yu (2020); Feng et al. (2022) extend this idea to multiple parties.
For the related case of VFL with small $|\mathcal{S}_{\mathrm{guests}}\cap\mathcal{S}_{\mathrm{labels}}|$, Sun et al. (2023) introduce few-shot Vertical Federated Learning, where a few epochs of training on the aligned dataset is followed by semi-supervised learning on the non-intersecting entities of each client’s dataset. To generate labels, guests cluster the gradients they receive from the aligned samples. While this enables remarkable model performance, there is unwanted computational overhead required to generate the labels. Moreover, this is a form of label inference attack! Not only is this a breach of privacy to some degree, this clustering could be used downstream by a malicious agent to perform feature inference attacks (see Section 2.4.3).
3.3 Greedy and Localized Learning
Algorithms which update parameters in hidden layers of ANNs based on local or auxiliary objectives (as opposed to the global label loss) have recently seen renewed interest. In the scope of VFL training, they enable some useful properties for ANN training.
The primary drawbacks of BP in the scope of VFL are the inherent locking synchronization constraints, described in Jaderberg et al. (2017). In order to update its weights, a layer must wait for the backward signal to reach from the output layer (backward locking). In order to begin updating any layer, a full forward pass must first propagate to the end of the network (update locking). In order for a layer to evaluate an input to begin with, a forward pass must propagate to it (forward locking). These strict locks at the split layer are exactly the failure points for VFL training– forward locking causes training to fail if a guest does not communicate activations, backward locking causes training to fail if the host does not communicate gradients.
By eschewing the wait to evaluate the global objective, most greedy algorithms are backward unlocked if not update unlocked. In order to partially address forward locking for model parallel training Belilovsky et al. (2020) propose a replay buffer mechanism. Each agent responsible for training a module uses a replay buffer to store activations from the previous module’s agent. During a forward pass, the agent reads an activation from the buffer. If the forward signal has propagated through the previous module, it would have written the latest activation to the buffer. If not, the previous activation is reused.
4 Proposed method
DVFL splits the training into two phases coordinated by the owner. The first is an unsupervised and decoupled greedy guest-host representation learning phase. In this phase, the guest and host models learn feature extractors. The second is a supervised owner transfer-learning phase where the label owner learns to aggregate encodings from the host space and produce label predictions.
Within the guest-host phase, participants train their models asynchronously on their own obejctives. This allows some flexibility regarding when participants communicate and when they train. Since guests do not rely on feedback from hosts to update their models, host or communication faults are irrelevant to the guest models. Hosts store activations from guests as they arrive and use the latest message for their forward pass, so a missing input on the forward pass is no longer critical.
After hosts and guests complete training, the label owner trains its model. For each entity in $\mathcal{S}_{\mathrm{labels}}$, the owner requests guests and hosts to send its encoded and aggregated features. The owner uses this as an input for its task model, e.g. a classifier.
We present a DVFL approach to training split neural networks in the follwing subsections. The resultant algorithm is formalized in Algorithm 1. We also show how DVFL guests might exploit samples outside the joint sample space $\mathcal{S}_{\mathrm{guests}}$ to train better guest and host models.
4.1 The DVFL Hierarchy
Guests: The role of a guest in DVFL is identical to a VFL guest’s (see Section 2.1), i.e. guests $g_{i}\in\mathcal{G}$ are the agents $g_{i}$ who own datasets $\mathcal{D}_{i}$ with records of features $\mathcal{F}_{i}$ for the set of entities whose indices are present in $\mathcal{S}_{i}$. Each guest has a guest model and an unuspervised objective.
Hosts: Hosts are members of the set $\mathcal{H}$ of agents $h_{i}$ which own aggregating host models, which accept the concatenation of all guest model outputs as input. Each $h_{i}\in\mathcal{H}$ has:
  1. 1.
    Host model $m_{h,i}(\text{ }\cdot\text{ };\theta_{h,i}):\mathbb{R}^{k_{1}}\times\mathbb{R}^{k_{2}}\times\dots\times\mathbb{R}^{k_{|\mathcal{G}|}}\to\mathbb{R}^{k_{h,i}}$ whose task is to aggregate and compress the data it receives from guests. It has its own unsupervised objective.
  2. 2.
    Single-reader single-writer “communication” registers $\mathbf{B}_{j,i}\forall j\in\{1,2,\dots,|\mathcal{G}|\}$. Each guest has a register on $h_{i}$, and writes its output to it. The host reads it to form its model input.
  3. 3.
    Input replay buffer $\mathbf{A}_{i}$. The host $h_{i}$ uses this to store a history of model inputs, if it wishes to reuse them.
Each host model has the same architecture but its own independent parameters $\theta_{h,i}$. It is possible that $\mathcal{H}\cap\mathcal{G}\neq\phi$.
Owner: The owner agent $o$ executes each training and communication round. It has access to ground truth labels, and is responsible for training a “final” or “owner” model on concatenated representations from members of $\mathcal{H}$ (i.e. it is a transfer learning model). The owner has the following:
  1. 1.
    Labels $L=\left\{\mathbf{y}_{j}\in\mathbb{R}^{\text{out}}:\text{label }\mathbf{y}_{j}\text{ of }\mathbf{x}_{j},j\in\mathcal{S}_{\mathrm{labels}}\right\}$:
    • $\mathcal{S}_{\mathrm{labels}}$ is the set of sample indices for which the host has the corresponding labels
    • We assume that the intersection of all the participants’ indices is not null to allow for training, i.e., $\left|\mathcal{S}_{\mathrm{guests}}\cap\mathcal{S}_{\mathrm{labels}}\right|>0$.
  2. 2.
    Final model $m_{o}(\text{ }\cdot\text{ };\theta_{o}):\mathbb{R}^{k_{H}\cdot{|\mathcal{H}|}}\to\mathbb{R}^{\text{out}}$ with parameters $\theta_{o}$.
The label owner may be a member of $H\cup G$. In any case, we refer to it shorthand as “the owner”. Figure 2 illiustrates a DVFL system with three guests and two hosts.
Figure 2: A DVFL system with three guests and two hosts. Guests train their local models on unsupervised objectives and hosts also train their aggregating models on unsupervised objectives. After that, the label owner trains a transfer learning model (such as a linear classifier head) on the encodings from the hosts.
4.2 Asynchronicity in DVFL
To avoid a single point of failure, host models should continue training in the absence of inputs from some guests. We adapt the use of activation replay buffers of Belilovsky et al. (2020) to facilitate training under guest and communication faults. We use single-reader single-writer “communication” registers so that hosts can reuse old activations from failed guests. A host reads the latest messages in its registers and uses them to fulfill its model inputs.
If hosts wish to train for more iterations than they receive activations for, they may store their model inputs in replay buffers. They may then reuse inputs as needed. We discuss these mechanisms in further detail in Appendix D.
The task of a host’s model is only to learn an encoding of features in the joint guest domain, which is not an instance-specific task. Therefore, a host may still update its model meaningfully with some out-of-date activations.
4.3 Data beyond the intersection
Until now, we have assumed the input datasets to each model are “aligned”, viz. they only contain records for members of $\mathcal{S}_{\mathrm{guests}}\cap\mathcal{S}_{\mathrm{labels}}$. The learning target of the guest models are to compress any data they receive as input, i.e. to encode features of a realization of guest domain’s data distribution. Similarly, each host learns to aggregate and compress encodings from all guests. These tasks depend on the input data’s domain (i.e. not on its instance). Therefore, guests may train their local models on features of $\mathbf{x}_{j}$ even when $j\in\mathcal{S}_{\mathrm{guests}}\setminus\mathcal{S}_{\mathrm{labels}}$.
4.4 DVFL for Split Training of Neural Networks
Given the DVFL hierarchy, ANNs are natural choices for participants’ models. Unsupervised neural networks are feature extractors Intrator (1992); Becker & Plumbley (1996), which is precisely the role of guest and host models. Similarly, the strong task performance of ANNs have been the primary driver of deep learning over the past few years, so they are suitable for owners, too. Indeed, just as SplitNN is a popular implementation of VFL, we expect ANNs to be common models for DVFL participants.
Figure 1 illustrates the relationship between DVFL and VFL in the context of a global ANN model. We may say $m_{i}$s are all ANNs with parameters $\theta_{i}$, $m_{h,j}$ are also ANNs whose input is the concatenation of the outputs of all $m_{i}$s, and $m_{o}$ is an ANN whose input is the concatenation of the outputs of all $m_{h,j}$s.
Algorithm 1 DVFL training for split neural networks
Input: Guest datasets $\mathcal{D}_{i}$, optimizers $\text{optim}_{g_{i}}$ $\forall i\in$$\{o\}\cup$$\{g_{1},\dots,g_{|\mathcal{G}|}\}\cup\{h_{1},\dots,h_{|\mathcal{H}|}\}$, communication schedule
Guest Training Round:
1:  for each guest $g_{i}$ in parallel do
2:     Take the next batch $\mathbf{x}_{t,\mathcal{F}_{i}}$ from dataset $\mathcal{D}_{i}$
3:     $\hat{\mathbf{x}}_{t,i}\leftarrow m_{i}(\mathbf{x}_{t,\mathcal{F}_{i}};\theta_{i})${Guest forward pass}
4:     if communication round then
5:        for each host $h_{j}\in\mathcal{H}$ do
6:           $\mathbf{B}_{i,j}\leftarrow\hat{\mathbf{x}}_{t,i}$
7:        end for
8:     end if
9:     Compute an unsupervised loss $\mathcal{L}_{i}(\hat{\mathbf{x}}_{t,i})$
10:     $\theta_{i}\leftarrow\text{optim}_{i}(\theta_{i},\nabla_{\theta_{i}}\mathcal{L}_{i})${Update guest model}
11:  end for
Host Training Round:
1:  for each host $h_{i}$ in parallel do
2:     if communication round then
3:        Read $\mathbf{B}_{j,i}\forall j\in\{1,2,\dots,|\mathcal{G}|\}$
4:        Concatenate all $\mathbf{B}_{j,i}$ and append to $\mathbf{A}_{i}$
5:     end if
6:     Read vector $\mathbf{b}_{t,i}$ from $\mathbf{A}_{i}$
7:     $\hat{\mathbf{b}}_{t,i}\leftarrow m_{h,i}(\mathbf{b}_{t,i};\theta_{h,i})${Host forward pass}
8:     Compute an unsupervised loss $\mathcal{L}_{h,i}\left(\hat{\mathbf{b}}_{t,i}\right)$
9:     $\theta_{h,i}\leftarrow\text{optim}_{h_{i}}(\theta_{h,i},\nabla_{\theta_{h,i}}\mathcal{L}_{h,i})${Update host model}
10:  end for
Owner executes:
1:  repeat
2:     Guest Training Round
3:     Host Training Round
4:  until convergence or a fixed number of iterations
5:  repeat
6:     for each entity $\mathbf{x_{j}}$, $j\in\mathcal{S}_{\mathrm{guests}}\cap\mathcal{S}_{\mathrm{labels}}$ do
7:        for all guests in parallel do
8:           Get features $\mathbf{x}_{j,\mathcal{F}_{i}}$ pertaining to entity $\mathbf{x_{j}}$
9:           Pass features through guest models and send activations to hosts
10:        end for
11:        for all hosts in parallel do
12:           Concatenate all received activations
13:           Pass activations through host models $m_{h,i}$ with parameters $\mathbf{\theta}_{h,i}$
14:           Send all activations to the owner
15:        end for
16:        Owner concatenates all inputs to form $\mathbf{x}_{j,\mathrm{enc}}$
17:        $\hat{\mathbf{y}}_{j}\leftarrow m_{o}({\mathbf{x}}_{j,\mathrm{enc}};\theta_{o})${Owner forward pass}
18:        Compute loss against labels $\mathcal{L}_{o}(\hat{\mathbf{y}}_{j},\mathbf{y}_{j})$
19:        $\theta_{i}\leftarrow\text{optim}_{o}(\theta_{o},\nabla_{\theta_{o}}\mathcal{L}_{o})${Update owner model}
20:     end for
21:  until convergence or a fixed number of iterations
5 Experiments
As a toy example, we consider a system of $|\mathcal{G}|$ guests which can each see $\frac{1}{|\mathcal{G}|}^{th}$ of an handwritten digit from the MNIST dataset. For our experiments, we set $|\mathcal{G}|=4$ i.e. $g_{1}$’s dataset contains the top $28\times 7$ pixel patch of an MNIST image, $g_{2}$’s dataset contains the second , $g_{3}$’s contains the third, and $g_{4}$’s contains the bottom $28\times 7$ patch (see Fig. 3).
Figure 3: For our experiments, the input received by each guest for a given sample, as well as the reconstruction its autoencoder model produces.
We use a learnable decoder head and MSE reconstruction as an unsupervised objective.
Baseline
DVFL is an alternate strategy to VFL. Naturally we compare DVFL for ANNs to SplitNN, a direct application of VFL to ANNs. Naive VFL implicitly employs the “wait” protocol discussed in Sec 2.4.1. There are some simple additions to SplitNN that would allow a more meaningful comparison. Firstly, if the host sets a deadline, we may continue training despite missing inputs by skipping (we call this SplitNN-skip) or by supplying replacement values. For replacement values, we try a zero tensor (SplitNN-zeros). Hosts may ask guests to send the identities of the entity to which each of their messages belong. This way, a late message cannot be believed to be intended for the next iteration. Results on more baselines and datasets are presented in Appendix A. The choice of baseline is discussed in Appendix C.
5.1 Fault Simulation
Six hyperparameters control the fault model described in Section 2.4.1. $R^{(d)}_{i}$ is the fault rate for a live $i\in\{\mathrm{connection},\mathrm{guest},\mathrm{host}\}$. $R^{(u)}_{i}$ is the chance a dead $i$ rejoins the network before the next request. We discuss how exactly the faults are simulated in Appendix B.2. We assume no faults during label supervision or inference.
5.2 Experimental Setup
Table 1: We compare DVFL to standard VFL. Here, we report the median and twice the standard deviation of MNIST test performance when trained with various entities $i\in\{\mathrm{connection,guest,host}\}$ susceptible to fail at a rate of $R^{(d)}_{i}=0.3$. A failed resource has $R^{(u)}_{i}=\mathbb{P}_{\mathrm{rejoin}}\in\{1,0.5,0.1\}$ to be available next time it is requested. SplitNN, the baseline VFL algorithm, cannot train at all when there are any faults. On the other hand, our DVFL implementation degrades gracefully with faults. Using an explicit fault handling strategy within VFL does allow it to train. But the degradation with faults is sharper than DVFL and the variance in model performance is higher. Under fault-free circumstances, SplitNN slightly outperforms our DVFL implementation. With minor adjustments, SplitNN can tolerate some faults, but our approach is more resilient and stable.
Strategy Test Accuracy (%) with 0.3 Loss Rate
No Faults Communication Loss Guest Loss Host Loss
$\mathbb{P}_{\mathrm{rejoin}}=$ N/A $\mathbb{P}_{\mathrm{rejoin}}=$ 1 $\mathbb{P}_{\mathrm{rejoin}}=$ 0.5 $\mathbb{P}_{\mathrm{rejoin}}=$ 0.1 $\mathbb{P}_{\mathrm{rejoin}}=$ 1 $\mathbb{P}_{\mathrm{rejoin}}=$ 0.5 $\mathbb{P}_{\mathrm{rejoin}}=$ 0.1 $\mathbb{P}_{\mathrm{rejoin}}=$ 1 $\mathbb{P}_{\mathrm{rejoin}}=$ 0.5 $\mathbb{P}_{\mathrm{rejoin}}=$ 0.1
DVFL-NN 97.30$\pm$0.08 97.24$\pm$0.10 97.19$\pm$0.14 97.03$\pm$0.17 97.11$\pm$0.11 97.08$\pm$0.16 96.87$\pm$0.20 97.27$\pm$0.12 97.24$\pm$0.14 97.02$\pm$0.33
VFL (SplitNN) 97.80$\pm$0.17 N/A N/A N/A N/A N/A N/A N/A N/A N/A
VFL (SplitNN-skip) 97.84$\pm$0.16 97.50$\pm$0.54 96.60$\pm$0.51 77.21$\pm$5.34 97.44$\pm$0.57 96.65$\pm$0.45 80.22$\pm$4.31 97.65$\pm$0.41 97.59$\pm$0.50 97.00$\pm$0.89
VFL (SplitNN-zeros) 97.79$\pm$0.15 97.64$\pm$0.38 97.53$\pm$0.40 96.89$\pm$0.36 97.68$\pm$0.40 97.59$\pm$0.40 96.45$\pm$0.50 97.68$\pm$0.30 97.65$\pm$0.37 97.01$\pm$0.93
Performance under Faults
In order to show that it is possible to use DVFL even with crash faults, we train DVFL with $|\mathcal{H}|=4$ hosts under $R^{(d)}_{i}=0.3,R^{(d)}_{j}=0$ where $i\in\{\mathrm{connection},\mathrm{guest},\mathrm{host},\phi\},j\neq i$. Given a $R^{(d)}_{i}\neq 0$, we try $R^{(d)}_{i}\in\{0.1,0.5,1.0\}$.
Effect of Redundancy
In order to show that redundancy in hosts improves model performance and fault tolerance, we simulate DVFL training with $|\mathcal{H}|$ = 1, 2, 3, and 4 hosts. We measure test accuracy given $R_{\mathrm{connection}}\in\{0,0.3,0.6\}$ during training.
Learning from a Limited Intersection
We first fix a small intersection size, where all members of the intersection are labeled. We generate the remaining dataset for each device by shuffling the out-of-intersection samples and taking windows such that the remaining sample space is distributed evenly across all four guests. For example, if we wish to use $1024$ labeled samples, we shuffle samples indexed $1024-59999$ and split the $58976$ samples evenly, such that each guest gets $14744$ unique MNIST patches + $1024$ intersecting MNIST patches = $15768$ total samples. We try this experiment with $128$, $256$, $512$ and $1024$ labeled samples. We compare with SplitNN trained on only the labeled samples.
6 Results and Discussion
[…]
Figure 4: Model performance improves with redundancy, especially when there is a nonzero crash rate.
6.1 Graceful Degradation with Faults
We observe (Table 1) that model performance degrades with faults, but not catastrophically (unlike VFL with SplitNN). In ideal conditions, DVFL performance is slightly poorer than SplitNN. This is not unexpected, since our models are shallow. Wang et al. (2021) show performance improves with the length of backpropagation in blockwise-trained networks, and suggest it is due to increasing mutual information between shallower and deeper layers. Further, given a model, the reconstruction loss guides the feature extractor to features that will be useful for reconstruction– not necessarily good for classification. It is expected that the gap between DVFL and SplitNN will reduce with more general loss functions van den Oord et al. (2019); Chen et al. (2020b); Caron et al. (2020); Zbontar et al. (2021) and with self-supervised classification objectives. Further, observe that the degradation with faults in DVFL is much more graceful than our SplitNN variants. Not only does the SplitNN-wait/skip/stale performance drop more sharply, the variance increases more than it does with DVFL.
Guest faults are the worst
In the case of DVFL, the most severe type of fault is the guest loss. For example, when $R^{(d)}_{\mathrm{guest}}=0.3,R^{(u)}_{\mathrm{guest}}=0.1$ the accuracy drops from 97.30% to 96.87% while with communication and host loss it is 97.03 and 97.02% respectively. The performance of the combined model depends on unique outputs from guests– there is no redundancy. With a host or communication fault, guests are still able to update their own models (and there might be redundant hosts, too). In the case of SplitNN-skip and SplitNN-host, a communication fault is just as bad– a guest cannot send forward passes to (or receive backward passes from) the single host whether it or the connection is down. A host fault is equivalent to skipping that round for training.
$R^{(u)}=1$ simulates single-message loss, since the next time a message is sent or received, the participant is sure to have rejoined. $R^{(u)}=0.5$ simulates short bursts of loss. $R^{(u)}=0.1$ simulates appreciable downtime.
Tolerating some faults pays off
Induced by the artificial input, the Dropout effect Srivastava et al. (2014) leads to a reduction in bias. This competes against the loss of information in the model. In the case of DVFL, models do not accept feedback, so the scope for information loss is smaller. In the case of SplitNN with missing input imputation, models do not respond to faults as well as DVFL. Still, the test accuracy degrades somwewhat gracefully for this reason.
Other failure modes
In this study, we only examine loss faults. Other typical failure modes for distributed systems include arbitrary delays, incorrect order, and Byzantine faults. For the forward pass, the first two might be mitigated within VFL itself. Since hosts request encoded features from a particular sample or batch of samples, they may ask guests to send back the identity of those samples along with the features (as we assumed in the variants of SplitNN. While the privacy is dubious, there is precedent in VFL: see Sun et al. (2021); Yang et al. (2023); Jiang et al. (2022). But, in order to correctly serialize, the host must maintain a cache of unused messages. For the backwad pass, if a guest has sent two outputs before receiving a single gradient, it has no way of knowing which output it corresponds to. DVFL is immune to these issues, since guests and hosts train independently to each other.
6.2 Redundancy Improves Performance
We observe (Figure 4) that test accuracy improves with the number of hosts with all other system parameters fixed. This is not surprising, as the owner’s aggregation of host results can be seen as a form of bagging Breiman (1996). As a corollary, the degradation of performance with communication faults is more graceful with increasing redundancy.
The role of the DVFL host
Observe that host faults do not hurt DVFL performance very much. DVFL hosts’ primary function is aggregation. A DVFL host can achieve this without any learnable parameters, by simply concatenating inputs and passing them onto the owner when requested. The other role a DVFL host might play is to project the concatenated host output to a smaller dimension. What the host model must learn is to preserve useful features in the new dimension.
Table 2: We exploit data outside the sample intersection to train DVFL’s guests and hosts. For the owner model, we use labeled entities common to all datasets. This performs better than VFL, which can only use the common samples.
Strategy Test Accuracy (%)
# of Labeled Samples
128 256 512 1024
DVFL 76.45 82.51 85.18 88.90
VFL (SplitNN) 65.07 77.51 83.09 87.26
6.3 Out-of-Intersection Training
Although VFL with SplitNN outperforms our DVFL ANN algorithm in our experiments with perfect conditions, DVFL significantly outperforms VFL on a limited labeled intersection. DVFL’s advantage is greater with fewer labels (Table 2).
6.4 Security & Privacy
DVFL achieves these results while also completely eliminating the possibility of gradient-based inference attacks. As reviewed in 3, complex mechanisms can be developed on top of SplitNN to provide privacy guarantees within some bounds. DVFL achieves complete gradient privacy without any overhead or model convergence issues.
7 Conclusion and Future Work
We present Decoupled Vertical Federated Learning, a strategy for joint learning on vertically partitioned data. Instead of adding complexity to cope with the data leakage and synchronization associated with BP, we eschew feedback altogether. Combined with the ability to reuse feedforward signals, this allows for fault tolerance and privacy in training. Training may continue even when there are faults, or there are a limited number of samples common to all guests.
DVFL also provides a high degree of privacy. Malicious or compromised guests can no longer use gradient feedback to figure out possibly sensitive information. Existing techniques to enforce privacy guarantees either hamper model convergence, are expensive to implement or both.
Our results suggest its performance to be comparable to SplitNN on various machine learning tasks. Unlike SplitNN, DVFL degrades gracefully with various kinds of loss faults. Even with explicit fault handling strategies, SplitNN’s performance drops off more sharply than DVFL’s. In general, DVFL’s asynchronicity provides a great deal of engineering flexibility. Some of it is discussed in the Appendix.
The next step for DVFL is to generalize it for more tasks and general architectures. VFL is readily compatible with a variety of machine learning models (including ensemble models). With a judicious choice of training objective for guest models, DVFL may also be extended past ANNs.
Another possible direction for DVFL is the introduction of more elaborate fault handling strategies. In this work, the aggregating “host” uses the the latest activation received from a “guest” client. This allows for a simple implementation using single-reader single-writer registers. We observe that this leads to meaningful representation learning– empirically, we only observe a 0.26% drop in accuracy on MNIST even when there is a 30% crash fault rate. In a bid to close this gap further, hosts might build a model of the distribution of its inputs. Hosts could then sample missing inputs from its posterior conditioned on successful messages. A natural first step would be to use a running mean as an estimator.
Appendix A Results on More Datasets, Comparison to FedVS
In order to bolster the comparison to SplitNN, we evaluate the performance of DVFL on some benchmark vertically partitioned datasets. They are Credit Card Yeh & hui Lien (2009), Parkinson’s Sakar et al. (2019), CalTech-7 and Handwritten Markelle Kelly . Note that CalTech-7 is a subset of 7 classes CalTech 101 Li et al. (2022): Face, Motorbikes, Dolla-Bill, Garfield, Snoopy, Stop-Sign and Windsor-Chair. Both Handwritten and CalTech have 6 distinct “views”– features of distinct domains, e.g. spatial, frequency (Fourier), Gabor filter coefficients, etc.
FedVS Li et al. (2023) is a state-of-the-art VFL algorithm which aims to tackle both privacy and fault tolerance. There is a round of secret shares so that each guest holds an encrypted version of a horizontal section of other guests’ data and model parameters. The shares are arranged in such a manner that if $\leq L$ (a fixed system parameter) guests don’t send their outputs at the forward pass, there is enough redundant information in the other guests’ forward passes to account for it. In order to maintain privacy despite this sharing of data, guests quantize their models and data and apply randomly mask. This means there is no longer enough information for a malicious guest to exactly reconstruct the data it receives from other guests. But, this obscurity is at the cost of model performance. Further, this algorithm is contingent on the guests using polynomial networks Livni et al. (2014a; b).
In the following experiment, we compare our DVFL algorithm, SplitNN and FedVS. Our VFL and DVFL models are of the same depth as FedVS, but we use a smaller width. We also only use population normalization (as opposed to online batch and layer normalization). Therefore, the non-FedVS models have much fewer parameters and are simpler to train. We also take the liberty of using non-polynomial activations (Leaky ReLU). Model details and hyperparameters are in Appendix B. $|\mathcal{H}|=1$ for these experiments.
Table 3: On a variety of typical vertically partitioned datasets, we see that DVFL performs comparably to SplitNN even under perfect conditions. Even on tabular datasets harder (especially for ANNs) than MNIST such as Credit Card and Parkinson’s, DVFL model performance is comparable to SplitNN. Despite using a smaller model, both outperform the results cited in the FedVS paper Li et al. (2023).
Strategy Test Accuracy (%)
Dataset
Credit Card Parkinson’s CalTech-7 Handwritten
DVFL 82.66 85.76 99.00 98.25
VFL (SplitNN) 84.11 86.64 99.66 98.25
FedVS Li et al. (2023) [^2] Results extracted from graphs. Raw figures were unavailable and the authors did not respond to our inquiries. 81.99 81.83 95.98 98.25
The feature splits and train-test splits for each dataset are identical to the FedVS paper Li et al. (2023). Further discussion about these results is available in Appendix C.
Appendix B Models, Hyperparameters and Implementation Details
B.1 Models
The “global”, single-participant neural network (see Figure 2) being split is a shallow multi-layer perceptron. It is parameterized by $W_{g}$ and $W_{h}$. $W_{g}$ is the dimension of the concatenated output of all guests viz. the input to a host. $W_{h}$ is the output dimension of a single host.
B.1.1 MNIST
The “guest” models have 2 layers, which squeeze the flattened $\frac{784}{|\mathcal{G}|}$-long input vector dimension to $\frac{400}{|\mathcal{G}|}$, and then to $\frac{W_{g}}{|\mathcal{G}|}$ with ReLU activation.
The “host” models squeeze the $W_{g}$-dimensional input to $\frac{W_{g}+3W_{h}}{4}$ and then to $W_{h}$, with Leaky ReLU activation (with negative half-plane slope = 0.01).
The “owner” model squeezes the $W_{h}\times|\mathcal{H}|$ dimensional input to $W_{h}$ and then to $40$ with a Leaky ReLU activation and finally to a $10$-class prediction with an implicit softmax via the cross entropy loss function.
For DVFL experiments, the guest, host, and owner modules are separate objects simulated to reside on separate devices. In order to produce a reconstruction, they also contain decoder MLPs which undo the squeezing operation. The guest decoder is activated by a sigmoid function (since the input image is normalized to the range $[0,1]$) and the host decoder is activated by a ReLU function (since the output of the guest model is also activated by a ReLU function and therefore cannot be negative).
For SplitNN experiments, the guest encoder architecture is identical to DVFL. The host model is a sequential module of the DVFL host encoder followed by the owner classifier.
We try DVFL for $W_{g}\in\{400,320,240,200\}$ and $W_{h}\in\{200,160,120\}$. The classification accuracies are presented in Table 4.
Table 4: Ablation study of model width (particularly, the dimensions of the outputs which are communicated) on model performance with DVFL. Our experiments in the main paper use $W_{g}=320$ and $W_{h}=160$.
$W_{h}$ Test Accuracy (%)
$W_{g}=200$ $W_{g}=240$ $W_{g}=320$ $W_{g}=400$
120 96.65 96.93 96.89 96.92
160 97.07 97.31 97.39 97.15
200 97.40 97.13 97.11 97.16
There two competing factors: the number of parameters (a monotonic function of $W_{g}\times W_{h}$) and the compression from host to guest. While more parameters implies a more generalizable (and hence hopefully accurate) model, an increase in $W_{g}$ for a fixed $W_{h}$ seems to suffer from some information loss, especially for larger $W_{h}$. It could be that the model eschews an encoding with more classification-relevant information when encodings with more reconstruction-relevant information are available thanks to the larger parameter space.
In the main paper, we use $W_{g}=320$ and $W_{h}=160$ for our experiments on MNIST. This choice enables good model performance while also demonstrating host models’ ability to compress the data.
B.1.2 VFL Datasets
Similar to MNIST, we use shallow multi-layer perceptrons for our results on vertically partitioned data in Appendix A. The model sizes for the guest $g_{i}$, host $h_{k}$ and owner $c$ are summarized in table 5, where $d_{i}$ is the input dimension of guest $g_{i}$.
Table 5: Model dimensions used in our experiments on tabular datasets. The dimension of the input to a guest $g_{i}$’s model is $d_{i}$. $W_{g}$ is the dimension of the concatenated output of the guests’ models. $W_{h}$ is the dimension of the output of one host’s model.
Dataset $W_{g}$ $W_{h}$ Activation
Guest Host
Credit Card $\sum_{i}\lceil\frac{3d_{i}}{4}\rceil$ 10 Leaky ReLU Leaky ReLU
Parkinsons $\sum_{i}\lfloor{\frac{d_{i}}{4}}\rfloor$ $\frac{\sum_{i}\lfloor{\frac{d_{i}}{4}}\rfloor}{4}$ Leaky ReLU ReLU
CalTech-7 $256\times|\mathcal{G}|$ $256$ Leaky ReLU ReLU
Handwritten $\sum_{i}\lceil\frac{3d_{i}}{4}\rceil$ $3\frac{\sum_{i}\lceil\frac{3d_{i}}{4}\rceil}{4}$ Leaky ReLU ReLU
B.2 Fault Simulation
We simulate faults based on six hyperparameters:
  1. 1.
    Communication Faults: During each communication round, for each pair $(g_{j},h_{i})$ we draw a sample from a uniform distribution with support $[0,1]$. We track the current status of each connection using flags.
    • $R^{(d)}_{\mathrm{connection}}$: If the connection between $(g_{j},h_{i})$ was alive on the last iteration, we kill it if the sampled value is greater than $R^{(u)}_{\mathrm{connection}}$.
    • $R^{(u)}_{\mathrm{connection}}$: If the connection between $(g_{j},h_{i})$ was dead on the last iteration, we revive it if the sampled value is lesser than $R^{(u)}_{\mathrm{connection}}$.
    In DVFL, we only write $g_{j}$’s activations to register $\mathbf{B}_{j,i}$ if. Similarly, for SplitNN, we check and update the flag before any forward pass from or backward pass to the split layer.
  2. 2.
    Guest Faults: Whenever a guest is called (i.e. there is a forward or backward pass), for each $g_{j}$ we draw a sample from a uniform distribution with support $[0,1]$.
    • $R^{(d)}_{\mathrm{guest}}$: If the guest $g_{j}$ was alive on the last call, we kill it if the sampled value is greater than $R^{(u)}_{\mathrm{guest}}$.
    • $R^{(u)}_{\mathrm{guest}}$: If the guest $g_{j}$ was dead on the last call, we revive it if the sampled value is lesser than $R^{(u)}_{\mathrm{guest}}$.
    In DVFL, we only update $\theta_{j}$ if the guest $g_{j}$ is alive. If it is a communication epoch, we only write $g_{j}$’s activations to $\mathbf{B}_{j,i}\forall i\in\{1,2,\dots,|\mathcal{H}|\}$ if the guest is alive. Similarly, in SplitNN, we only allow a guest to send a forward pass and receive gradients if it is alive.
  3. 3.
    Host Faults: During each training iteration, for each $h_{i}$ we draw a sample from a uniform distribution with support $[0,1]$.
    • $R^{(d)}_{\mathrm{host}}$: If the guest $h_{j}$ was alive on the last call, we kill it if the sampled value is greater than $R^{(u)}_{\mathrm{host}}$.
    • $R^{(u)}_{\mathrm{host}}$: If the guest $h_{j}$ was dead on the last call, we revive it if the sampled value is lesser than $R^{(u)}_{\mathrm{host}}$.
    In DVFL, we only update $\theta_{H,i}$ if the host is alive. If it is a communication epoch, we also drop the last item in input buffer $\mathbf{A}_{i}$. In the case of SplitNN, we assume the host is able to call on the guests before failing. When a SplitNN host fails, it does not compute any gradients.
B.3 Other Hyperparameters
For all experiments, we perform a hyperparameter search. The values we used for the results presented in this paper are below.
B.3.1 MNIST
For guests and DVFL hosts, we use an Adam optimizer with the default $\beta_{1}=0.9,\beta_{2}=0.999$. We use a learning rate of $10^{-3}$ and a weight decay of $10^{-5}$. For DVFL owners and SplitNN guests we use stochastic gradient descent with a learning rate of $10^{-2}$ and momentum of $0.5$. In general, guests were trained for 20 epochs, hosts for 40 and owners for 60. For SplitNN, the whole model was trained end-to-end with 60 epochs. For the experiments on limited intersecting datasets, owners/SplitNN hosts are trained for 160 epochs. In all cases, DVFL owner models and SplitNN host models were trained with early stopping.
B.3.2 VFL Datasets
Hyperparameters are presented in Table 6.
Table 6: Hyperparameters for the results presented in Appendix A.
Dataset Learning Rate
DVFL SplitNN
Guest Host Owner Guest Host
(Adam, $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.999) (Adam, $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.999) (SGD, momentum = 0.5) (Adam, $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.999) (Adam, $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.999)
Handwritten 2e-3 1e-3 3e-3 1e-3 1e-4
CalTech-7 5e-4 1e-2 3e-5 1e-4 1e-4
Parkinson’s 5e-3 4e-3 5e-4 1e-4 5e-5
Credit Card 1e-3 1e-3 1e-3 1e-3 1e-3
B.4 Implementation Details
We simulate parallel training by time multiplexing and using preallocated replay and communication tensors. The experiments were run on a variety of hardware, including mobile CPUs (Apple M2) and workstation GPUs (on a single NVIDIA RTX A5000 and single-GPU, and on a single NVIDIA Tesla K40).
Appendix C Choice of Baseline
C.1 SplitNN as a Baseline
In this work we present DVFL, an alternate strategy to VFL. We demonstrate its prowess through a DVFL implementation for ANNs. Since DVFL is altogether a new strategy to train on vertically partitioned data, it makes sense to benchmark it against the current status quo for VFL with ANNs, which is SplitNN Gupta & Raskar (2018); Ceballos et al. (2020).
SplitNN with a timeout
In its base form, the SplitNN algorithm does not have any explicit fault handling strategy. Implicitly (assuming that the coordination between hosts and guests is interrupt-driven), the hosts use the “wait” strategy. Training stops completely if inputs are missing. We also consider strategies using “timeout”. Timeout allows training to continue, although hampered. We handle timeouts by skipping or replacing the values. If a message appears after the deadline, it needs to be discarded lest it appear as if it were intended for the next iteration. This artificial baseline represents a best-case fault-tolerance for vanilla VFL. This way, it gets a fair shot against DVFL. Indeed, since DVFL suffers from some information loss (since feedback is very limited), VFL equipped with these strategies may still outperform DVFL at low fault rates.
Busy-waiting SplitNN
Under “wait”, training can be made to restart if the failed participant comes back online by busy-waiting viz. the repeatedly polling a failed guest. If the guests send sample IDs along with their features (and hosts cache messages until they are used), they may figure out which iteration a particular message is intended for. This way, no information is lost. However, until that failed participant is back online, training is held up. This is also a poor strategy from a systems perspective, since it wastes bandwidth and computational resources, especially when there is a message authentication in place. If a guest is laggy and the host polls it many times before it repsonds, the cache of messages the host must store will be very large. Neither in this case nor the timeout case is there a well-defined strategy to handle backward-pass faults.
C.2 Other Possible Baselines
DVFL addresses various VFL issues such as fault tolerance, privacy and sample size with one fell swoop. There are various mechanisms built upon standard VFL algorithms (including SplitNN) which aim to address some or the other of these issues. FedVS Li et al. (2023) is a work that addresses multiple, i.e. fault tolerance and privacy. FedVS might be a reasonable benchmark for DVFL, and we have presented some results against its paper in Appendix A. DVFL outperforms FedVS in perfect conditions. In the main paper, we showed that DVFL performance does not degrade much with faults with our generalized fault model. On the other hand, FedVS performance is completely fault tolerant, except under a much more stringent fault model. Beyond this, FedVS constrains guest models to use a specific type of model– polynomial networks Livni et al. (2014a; b). Moreover, FedVS only deals with faults during the forward pass. FedVS considers a restricted fault model: upto a fixed number of guests can be late to/fail to communicate their activations with the hosts. These are also true for most of its predecessors e.g. Chen et al. (2020a); Hu et al. (2019); Shi et al. (2022). DVFL does not require any such system parameter, and is tolerant to a broader class of faults where any participant can fail at any time, including during backpropagation.
Appendix D Communication and Space Complexity
D.1 Periodic Communication
The owner initiates every round of training as well as communication for all agents. Since the training of guests and hosts are decoupled, the owner may ask guests to communicate their latent activations (and therefore start host training rounds) at arbitrary intervals. In VFL, guest model updates are dependent on host feedback. So there is every-shot communication, i.e. guests send every single output to the host.
With one-shot communication such as FedOnce Wu et al. (2022), the volume of communication is greatly reduced, at the cost of model performance. Equivalently, Belilovsky et al. (2020); Siddiqui et al. (2023) demonstrate that updating all layers on local objectives on each forward pass performs better than fully training and freezing a layer before moving on to the next.
We study the tradeoff between communication cost and model performance by introducing a hyperparameter $K$, which is the period (in guest epochs) at which communication epochs occur. During a communication epoch, every iteration of the guest model iteration is followed by a communication round i.e. guests write their activations to the communication registers, and hosts read them, concatenate them and store the concatenations in their input buffers.
If there is a host training iteration at every guest training iteration communication round, after the guests complete $N$ epochs, the hosts will have
  1. 1.
    A model trained for $\left\lfloor\frac{N}{K}\right\rfloor$ epochs
  2. 2.
    The history of model inputs from each iteration in $\left\lfloor\frac{N}{K}\right\rfloor$ epochs of training
If the host must complete a certain number $M>\frac{|\mathcal{S}_{\mathrm{guests}}|}{\text{batch size}}\times\left\lfloor\frac{N}{K}\right\rfloor$ iterations of training, it reuses these activations until the required number of iterations is met.
We simulate VFL training with $4$ hosts and 00 fault rate. We measure test accuracy for the following communication periods:
  • Every guest epoch ($K=1$)
  • Every 5 guest epochs ($K=5$)
  • Every 10 guest epochs ($K=10$)
  • Only once, after 20 guest epochs ($K=20$).
Note that the dimension of the latent representations communicated by each guest to reach host is $80$, and each dimension corresponds to a 32-bit floating point number. The MNIST training dataset has $60000$ samples. If the total number of guest training epochs is $N$ and the period of communication epochs is $K$, the total communication cost per guest is ||×$|\mathcal{H}|\times\lfloor\frac{N}{K}\rfloor\times 60000\times 80\times 32$ bits. Each message received is saved by the host in the activation replay [^3] Although this is not necessary if the host will not reuse activations, viz. a host wants to train for $k$ iterations and it receives activations of $k$ iterations from the guests. This is discussed in more detail in Appendix D.2.
0055551010101015151515202020202525252596.996.996.996.99797979797.197.197.197.197.297.297.297.297.397.397.397.397.497.497.497.4Number of HostsMNIST Test Accuracy (%)
Figure 5: Model performance degrades with an increase in communication period, i.e. more communication is correlated with better performance.
Consistent with results from Belilovsky et al. (2020); Wu et al. (2022); Siddiqui et al. (2023), training the depth-wise split model parallely yields better model performance. We observe the tradeoff between communication cost and model performance. With 1.536GB, 0.307GB, 0.154GB and 0.077GB outgoing from each guest, the accuracies on the MNIST test set are $97.30\%$, $97.14\%$ $97.08\%$ and $96.84\%$ respectively.
D.2 Activation Replay Mechanism & its Computational Overhead
A host reads its model inputs from its activation replay buffer. Thanks to this, it can train asynchronously to the guests. The host can train independently to the guests as long as the guests have communicated at some prior point of time. In fact, that is the case in our experiments– we simulate host models’ training offline, after guest training. As discussed in Appendix D.1, this allows control over when the guests communicate.
Another way to look at this is: hosts do not need to train for the same number of iterations as guests communicate messages for. This includes the case where guests communicate every iteration, but hosts want to train more (or less) iterations. As we detailed in Appendix B.3, we actually implement this for our experiments. While we use 40 epochs to train the entire SplitNN model, we only use 20 epochs to train the DVFL guests. Guests need not remain online for any more time than required to train 20 epochs. So, not only does this allow for more flexibility, it also helps fault tolerance.
Computational Resources
In order to maintain this asynchronicity, hosts need to maintain the activation replay buffer. For each message it receives, the size of the buffer grows by batch $\text{batch size}\times W_{g}$. Further, these expensive memory read/write operations also add time complexity to DVFL hosts’ training. This overhead does not exist in standard VFL. This it is DVFL’s primary drawback [^4] The other source of overhead is the calculation of the self-supervised/unsupervised loss. For example, consider our experiments with MSE as the guest and host objectives. Along with our model, we need to train a decoder head that projects the guest output onto a space of the same dimension as input. Modern contrastive self-supervised losses e.g. Chen et al. (2020b); Zbontar et al. (2021) typically use a shallow projector network from the desired representation onto a latent space. The loss is calculated using the projected vectors. Intuitively, in the absence of a pre-computed ground truth label, the model must perform extra computation to generate the “self-”supervisory signals., although it can be mitigated to some extent since some models can be trained for fewer epochs as discussed above. However, it is important to note that it is not strictly necessary to save all activations. If a host trains for $\leq$ the number of epochs that guests train/communicate for, it does not need to reuse activations. In that case, a host model input can be deleted after it is used once.
[…]

### v2

@@
 Abstract
-Vertical Federated Learning (VFL) is an emergent distributed machine learning paradigm wherein owners of disjoint features of a common set of entities collaborate to learn a global model without sharing data.
-In VFL, a host client owns data labels for each entity and learns a final representation based on intermediate local representations from all guest clients.
-Therefore, the host is a single point of failure and label feedback can be used by malicious guest clients to infer private features.
-Requiring all participants to remain active and trustworthy throughout the entire training process is generally impractical and altogether infeasible outside of controlled environments.
-We propose Decoupled VFL (DVFL), a blockwise learning approach to VFL.
-By training each model on its own objective, DVFL allows for decentralized aggregation and isolation between feature learning and label supervision.
-With these properties, DVFL is fault tolerant and secure.
-We implement DVFL to train split neural networks and show that model performance is comparable to VFL on a variety of classification datasets.
+Vertical Federated Learning (VFL) is an emergent distributed machine learning paradigm for collaborative learning between clients who have disjoint features of common entities.
+However, standard VFL lacks fault tolerance, with each participant and connection being a single point of failure.
+Prior attempts to induce fault tolerance in VFL focus on the scenario of “straggling clients”, usually entailing that all messages eventually arrive or that there is an upper bound on the number of late messages.
+To handle the more general problem of arbitrary crashes, we propose Decoupled VFL (DVFL).
+To handle training with faults, DVFL decouples training between communication rounds using local unsupervised objectives.
+By further decoupling label supervision from aggregation, DVFL also enables redundant aggregators.
+As secondary benefits, DVFL can enhance data efficiency and provides immunity against gradient-based attacks.
+In this work, we implement DVFL for split neural networks with a self-supervised autoencoder loss.
+When there are faults, DVFL outperforms the best VFL-based alternative (97.58% vs 96.95% on an MNIST task).
+Even under perfect conditions, performance is comparable.
+footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
+Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
 1 Introduction
-Federated Learning McMahan et al.
-(2017), or FL, was introduced by Google researchers as a strategy for distributed learning, addressing communication efficiency and data privacy.
+Federated Learning (McMahan et al., 2017), or FL, was introduced by Google researchers as a strategy for distributed learning, addressing communication efficiency and data privacy.
 Distributed participants in FL training do not expose their data to any other party.
@@
 Among FL’s developments is Vertical Federated Learning (VFL).
-Standard FL can be said to have a “horizontal” or “sample-parallel” division of data, i.e.
-each participant holds unique samples within a shared feature space.
-On the other hand, VFL participants hold unique features of a common sample space, which is a “vertical” division, and is “feature-parallel”.
-This allows guest agents who individually have incomplete information about their target to learn meaningful joint representations without sharing data.
-A typical example: an imaging center, pathology lab and general hospital may come together to predict rare types of cancer, even though health regulations such as HIPAA may disallow them from collecting each others’ data.
-VFL guest agents undertake an “entity alignment” phase before training, where they ensure their datasets’ horizontal indexing is identical, i.e., all participants process features belonging to the same entity in any round of training.
-At each round, each guest passes its features to its local “guest model.” Guests share their local representations with a host.
-For example, a guest training a vertically federated ANN may send its activations.
-A boosted decision tree guest may send a weak learner.
-The host owns data labels and is aware of the identity of the sample or batch that the guests are processing at any time.
-The host also possesses a “host model” capable of aggregating these representations and making predictions against the labels.
-The host calculates a loss based on its aggregation and provides feedback (e.g.
-gradients) to the guests, which then begin the next round of training.
-We refer to the combined system of guest and host models as VFL models.
-VFL has found increasing success in recent years, with various real-world implementations and proposals Liu et al.
-(2023).
-However, VFL suffers from some flaws which make its implementation impractical where participants are few in number and each computationally powerful (as in cross-silo scenarios), and entirely infeasible otherwise.
-One such flaw is that the link between the guests and the host is a single point of failure.
-Training a batch of data is contingent on the exchange of intermediate results between the guests and the host.
-If the host waits indefinitely for all guests to communicate, a guest crashing or losing connection would completely halt training.
-If the host sets a deadline, some inputs to its model will remain unfulfilled.
-(In the latter case, a late message will also cause a catastrophic fault).
-If gradients from the host to the guests are lost or arrive in the incorrect order, guest model convergence is hampered.
-One might dismiss such events as infrequent or at least controllable in cross-silo settings.
-However, that is naive.
-When some of the guests (and even the host) could be geographically sparse and/or low-power devices, it is catastrophic.
-A fundamental limitation of VFL is imposed by the size of the joint sample space.
-For instance, databases in groups of large nationwide financial institutions may share many but not all customers.
-Guests must discard data outside the intersection, which could have been used to learn better guest models.
-In systems where the intersection is small, there may not be enough samples to train a VFL model altogether.
-In general, this also hinders VFL’s applicability to systems with large-scale participation.
-Consider again the hospital example.
-Suppose the task is now to predict bone marrow cancer.
-Along with the pathology lab, the imaging center and the general hospital, we also include data collected by an orthopaedic hospital.
-Even if the overlap of patients at the other facilities is significant, their overlap with patients of the orthopaedic hospital may be too small.
-Another drawback of VFL (and FL in general) is the risk of inference attacks from curious guests.
-Intuitively, a host makes predictions based on information it receives from all guests.
-The updates or feedback it returns is derived from this prediction and the ground truth.
-Therefore, feedback sent to one guest contains some mutual information with respect to the others and to the label.
-Luo et al.
-(2021); Fu et al.
-(2022) show that a malicious agent can engineer its messages to the host such that the label and the other guests’ features can be conveniently extracted from the gradients it receives in return.
-Attempts to safeguard VFL against inference attacks are generally based on differential privacy, homomorphic encryption, and multi-party techniques such as audits and shared secrets Yang et al.
+Standard FL can be said to have a “horizontal” or “sample-parallel” division of data: each participant holds unique samples within a shared feature space.
+Conversely, VFL participants hold unique features of a common sample space, which is a “vertical”, “feature-parallel” division.
+This allows guest agents who individually have incomplete and information about their target to learn meaningful joint representations without sharing data.
+VFL has received nascent interest, with some real-world proposals and applications (Liu et al., 2023).
+However, VFL requires careful synchronization which makes it difficult to engineer and scale.
+Critically, training is contingent on the exchange of intermediate results between the guests and the host.
+A crashed connection or participant on the backward pass means a missed model update.
+On the forward pass, it is catastrophic: an unfulfilled model input.
+We propose Decoupled Vertical Federated Learning (DVFL), where guests and hosts each train asynchronously on their own unsupervised objective before the label owner learns a transfer learning model from its labels.
+At an algorithmic level, DVFL avoids the forward and backward locking between hosts and guests.
+Hence, it is free of a single point of failure.
+Such “localized” training means that other participants can continue training even when one fails.
+Further, separating label inference from aggregation allows for redundant aggregators, further localizing faults.
+The exchange of intermediate results is also the root of many other limitations of VFL including inference attacks.
+Feedback received from the host has mutual information with labels and other guests’ features, which can be exploited by curious guests (mitigating this is typically costly in computation and performance: see Yang et al.
 (2023); Khan et al.
-(2023).
-These operations incur a high computational cost, which might not be feasible outside some cross-silo scenarios.
-Further, these methods hamper model convergence.
-Building intricate mechanisms to mitigate these issues inflates the complexity of the problem.
-In this work, we attack the source directly.
-Backpropagating gradients to guests allows inference attacks.
-The wait for a response or feedback causes synchronization issues.
-What if we eschewed end-to-end backpropagation (BP) altogether?
-We present Decoupled Vertical Federated Learning (DVFL), a novel strategy for ANNs to address these shortcomings.
-The underlying principle is blockwise learning, where gradient-isolated groups of layers of an ANN are trained on local objectives.
-Such greedily trained ANNs have been shown to perform comparably to those trained by BP Belilovsky et al.
-(2019); Journé et al.
-(2023).
-We decouple guest models from host models by training them on their own local objectives.
-Similarly, we decouple supervision from aggregation by training host models on local objectives.
-We use a final transfer learning model to make predictions against labels.
-This yields two useful system properties.
-One, this allows redundant hosts.
-Standard VFL would require hosts to share labels with each other, which is forbidden.
-Two, we may use separate datasets for feature extraction and label inference.
-With redundancy, even if a guest’s connection to one host fails, training may continue as normal on others.
-If hosts may fail, the redundancy also affords some degree of tolerance.
-By using separate datasets for inference and representation learning, we may train guest and host models on data which is not necessarily labeled and not necessarily a part of a joint sample space.
-This makes DVFL applicable to systems with small intersections of sample spaces.
-Further, the label owner’s supervised model can be trained completely offline, without the risk of faults.
-In Section 2, we provide a formal description of VFL and some algorithms which implement it for various machine learning models.
-We also elaborate on the shortcomings we hope to address with DVFL.
-In Section 3, we review some closely-related lines of research and use them to motivate DVFL.
-In Section 4, we provide a formal description of DVFL with an implementation for ANNs.
-We show some baseline results on benchmark machine learning tasks in Section 5.
-In the Appendix, we present more results and discuss our design choices.
+(2023)).
+Another is the inability to learn from features pertaining to samples outside the intersection of all participants’ sample spaces.
+DVFL does not suffer from these problems.
+Altogether, we expect DVFL to enable learning from vertically partitioned data in scenarios where they were so far infeasible, and help scale implementations where they weren’t.
 The main contributions of this work are:
   1. 1.
-    We present DVFL, a strategy which eschews host-to-guest feedback as well as label-to-host feedback and therefore has inbuilt redundancy, security and fault handling (including faults during the backward pass) while performing comparably to VFL.
+    We present DVFL, a strategy that eschews both host-to-guest and label-to-host feedbacks, and therefore enables redundancy, has inbuilt fault handling (including those during the backward pass) and inherent gradient privacy, and performs comparably to VFL.
   2. 2.
@@
 2 Background: Standard VFL
-Entities $\mathbf{x}$ are realizations of a fixed distribution $X$ (say).
+2.1 Notation
 The rows of a dataset are features of a unique entity $\mathbf{x}$.
-We may serialize these entities using a sample index.
-Let us say the set of sample indices corresponding to entities the host has labels for is $\mathcal{S}_{\mathrm{labels}}$.
-Let us say the entity with index $j$ is $\mathbf{x}_{j}$.
+We may serialize these entities using a sample index e.g.
+as $\mathbf{x}_{j}$.
+$g_{i}$ is the $i^{\text{th}}$ data-owning agent, member of set of guests $\mathcal{G}$.
+$\mathcal{F}_{i}$ is the set of features that guest $g_{i}$ records.
+$\mathbf{x}_{j,\mathcal{F}_{i}}$ are the features $\mathcal{F}_{i}$ of the $j^{\text{th}}$ entity.
+$m_{i}(\text{ }\cdot\text{ };\theta_{i})$ is $g_{i}$’s model, $\mathbb{R}^{|\mathcal{F}_{i}|}\to\mathbb{R}^{k_{i}}$ function parameterized by $\theta_{i}$, which is learnable.
+$\mathcal{S}_{i}$ is the set of sample indices for which $g_{i}$ has records.
+$\mathcal{S}_{\text{guests}}$ = $\bigcap_{i=1}^{|\mathcal{G}|}$.
+$h$ is the host agent, which owns labels.
+$\mathcal{S}_{\text{labels}}$ is the set of sample indices corresponding to labeled entities.
+$m_{h}(\text{ }\cdot\text{ };\theta_{h})$ is the host’s $\mathbb{R}^{k_{1}}\times\mathbb{R}^{k_{2}}\dots\times\mathbb{R}^{k_{|\mathcal{G}|}}\to\mathbb{R}^{\mathrm{out}}$ learnable model, which accepts the outputs of $m_{i}$ and makes a task-relevant prediction.
 2.2 The VFL Hierarchy
-Guests: $\mathcal{G}$ is the set of feature-owning collaborators: the guests. ||$\cdot$|| ⋅ | denotes cardinality. Each $g_{i}\in\mathcal{G}$ has its own
-  1. 1.
-    Dataset $\mathcal{D}_{i}=\left\{\mathbf{x}_{j,\mathcal{F}_{i}}\in\mathbb{R}^{|\mathcal{F}_{i}|}:j\in\mathcal{S}_{i}\right\}$, where $\mathcal{F}_{i}\subseteq\{1,2,\cdots,D\}$ is the subset of $D$ features that the guest owns, and $\mathcal{S}_{i}$ is the set of entities $g_{i}$’s dataset has records for.
-$\mathbf{x}_{j,\mathcal{F}_{i}}$ are the features $\mathcal{F}_{i}$ of $\mathbf{x}_{j}$.
-  2. 2.
-    Guest model $m_{i}(\text{ }\cdot\text{ };\theta_{i}):\mathbb{R}^{|\mathcal{F}_{i}|}\to\mathbb{R}^{k_{i}}$ with parameters $\theta_{i}$.
-The guest model $m_{i}$’s task is to concisely represent the features $\mathcal{F}_{i}$ of $\mathbf{x}_{j}$, for all sample indices $j$ that are in the intersection of all guests, i.e.
-$\forall j\in\mathcal{S}_{\mathrm{guests}}$, where $\mathcal{S}_{\mathrm{guests}}:=\bigcap_{i=1}^{|\mathcal{G}|}\mathcal{S}_{i}$ and pass them to the host $h$.
-Guests do not have access to other guests’ model parameters or data features.
-In general, they do not have access to labels, either.
+Guest: The guest role does not entail access to labels or other guests’ data.
+The guest’s private model $m_{i}$’s task is to concisely represent the features $\mathcal{F}_{i}$ of $\mathbf{x}_{j}$ $\forall$ sample indices $j\in\mathcal{S}_{\text{guests}}\cap\mathcal{S}_{\text{labels}}$.
+Guests obtain these encodings and pass them to the host $h$.
 Host: The host agent $h$ executes every round of training.
@@
 The host, given representations for all available features of $\mathbf{x}_{j}$, may use its own model to make useful predictions for $\mathbf{x}_{j}$.
-The host has
-  1. 1.
-    Labels $L=\left\{\mathbf{y}_{j}\in\mathbb{R}^{\text{out}}:j\in\mathcal{S}_{\mathrm{labels}}\right\}$:
-    • $\mathcal{S}_{\mathrm{labels}}$ is the set of sample indices for which the host has the corresponding labels
-    • We assume that the intersection of all the participants’ indices is not null to allow for training, i.e., $\left|\mathcal{S}_{\mathrm{guests}}\cap\mathcal{S}_{\mathrm{labels}}\right|>0$.
-  2. 2.
-    Host model $m_{h}(\text{ }\cdot\text{ };\theta_{h}):\mathbb{R}^{k_{1}}\times\mathbb{R}^{k_{2}}\times\dots\mathbb{R}^{k_{|\mathcal{G}|}}\to\mathbb{R}^{\text{out}}$ with parameters $\theta_{h}$.
-It is possible that $h\in\mathcal{G}$ i.e. the host may also be one of the guests. In such situations, the host is often known as the “active party” and guests as “passive parties.”
 If the host has access to the full computational graph, it may adjust all parameters to better predict label $\mathbf{y}$, usually by minimizing the expected value of some loss function.
@@
 where $\hat{\mathbf{x}}_{j,i}\equiv m_{i}(\mathbf{x}_{j,\mathcal{F}_{i}};\theta_{i})\in\mathbb{R}^{k_{i}}$.
-2.2 Private Set Intersection & Entity Alignment
-All guests process entities in the same order, each passing the next entity to its local model.
-For this to happen, all guests must know the set of common entities $\mathcal{S}_{\mathrm{guests}}\cap\mathcal{S}_{\mathrm{labels}}$.
-This is usually achieved using entity resolution (if required) and some form of Private Set Intersection (PSI).
-PSI is a multiparty computation of the intersection of sets belonging to each party without revealing members outside the intersection, typically by exchanging encrypted versions of the set Freedman et al.
-(2004).
-Entity resolution is a multiparty computation to generate a set of unique identifiers from features which are not strictly unique or precise.
-Entity resolution may be performed by comparing cryptographic long keys Hardy et al.
-(2017); Nock et al.
-(2018).
-After the host obtains the results of PSI, there is a phase of entity alignment.
-In this phase, the host decides the order of entities which all guests must follow and informs them.
+It is possible that the host may also be one of the guests. In such situations, the host is often known as the “active party” and guests as “passive parties.”
 2.3 Split Training for Neural Networks
-A faithful implementation of VFL for ANNs is SplitNN Gupta & Raskar (2018).
-SplitNN was originally proposed as an algorithm to distribute the training of a global neural network across two agents by splitting the network depth-wise into two.
+SplitNN (Gupta & Raskar, 2018) is a faithful implementation of VFL for NNs.
+It was originally proposed as an algorithm to distribute the training of a global neural network across two agents by splitting the network depth-wise into two.
 At the “split layer”, forward and backward signals are communicated between the two devices.
 [^1] One might say that the standard SplitNN addresses a special case of VFL where the host has access only to labels and the (single) guest has access only to features.
-The original paper Gupta & Raskar (2018) also presents a protocol for training when the features are horizontally distributed across $N$ guests.
+The original paper (Gupta & Raskar, 2018) also presents a protocol for training when the features are horizontally distributed across $N$ guests.
 Ceballos et al.
@@
 Figure 1: Diagram illustrating the distributed training of a neural network under SplitNN, VFL with SplitNN, and DVFL. The two input data for VFL and DVFL are partial features of the same entity.
-Effectively, the guest models $m_{i}$ are ANNs and host model $m_{h}$ consists of a concatenation layer $m_{h}^{0}$ followed by an ANN $m_{H}^{1}$.
+Effectively, the guest models $m_{i}$ are NNs and host model $m_{h}$ consists of a concatenation layer $m_{h}^{0}$ followed by an NN $m_{h}^{1}$.
 2.4 Impracticalities in VFL Training
-Vertical Federated Learning is contingent on implicit assumptions regarding the reliability, trustworthiness and scale of participation.
-Subversion of these assumptions can be catastrophic.
+VFL is contingent on implicit assumptions on the reliability, trustworthiness and scale of participation.
+As discussed in the introduction, subversion of these assumptions can be catastrophic.
 A lost guest output or gradient is a single point of failure.
@@
 And, it is possible for a malicious guest to infer private data from feedback.
-Thus far, the application of VFL has been limited to a small number of cross-silo settings where the silo owners are powerful enough to assume reliability and to afford the overhead of privacy guarantees.
-Further, cross-silo participation is small in number so scalability is not a concern.
-VFL is altogether infeasible for a large class of systems where devices are many or have low compute and communication bandwidths.
+We further analyze these shortcomings here.
 2.4.1 Single Points of Failure
@@
 We consider the following fault model for a round of training:
-  • Guest faults: A guest device may be unable to compute its output, update its model, or communicate.
+  • Guest faults: A guest may be unable to compute its output, update its model, or communicate.
   • Host faults: The host itself may fail, and be unable to receive inputs or calculate gradients.
   • Connection faults: The communication link between a guest and the host may be dropped during the forward or backward pass.
-From a VFL host’s perspective, a guest fault and communication fault manifest as a missing input to its model $m_{h}(\text{ }\cdot\text{ };\theta_{h}):\mathbb{R}^{k_{1}}\times\mathbb{R}^{k_{2}}\times\dots\mathbb{R}^{k_{|\mathcal{G}|}}\to\mathbb{R}^{\text{out}}$ during training.
+From a VFL host’s perspective, a guest fault and connection fault manifest as a missing input to its model $m_{h}(\text{ }\cdot\text{ };\theta_{h})$ during training.
 The host may employ one of two strategies:
@@
     Wait: The host may wait until all input arrives.
-    • Under a guest fault or communication fault, the host will remain waiting indefinitely, and training will not continue.
+    • Under a guest fault or connection fault, the host will remain waiting indefinitely, and training will not continue.
     • If the host polls an unresponsive guest, it may receive an input intended for the next iteration, which will lead to incorrect model evaluation.
@@
     Timeout: The host may set a deadline for each round.
-    • Under a guest fault or communication fault, training will fail since the host model is missing input.
+    • Under a guest fault or connection fault, training will fail since the host model is missing input.
     • If a message is successful but late, the model is still missing input.
     • If the model skips that training round and moves onto the next, it may receive the late message from the previous round, which will lead to incorrect model evaluation.
-Therefore, a communication or guest fault leads to a catastrophic failure for training.
+Therefore, a connection or guest fault leads to a catastrophic failure for training.
 When the host fails, the guest models are unable to update their parameters.
@@
 Fu et al.
-(2022) demonstrate that it is possible for a malicious agent to directly infer the true class from gradients (e.g.
+(2022a) demonstrate that it is possible for a malicious agent to directly infer the true class from gradients (e.g.
 by analyzing their sign) for many common loss functions during training.
-The authors also show that after VFL training is complete, it is possible for a curious guest to train its own prediction model if it can gain access to a very small number of labels.
-If a malicious guest inflates its learning rate during training, the aggregator gives more importance to results from its model.
-This way, a prediction model trained only on the features of the malicious guest is almost identical to the aggregator model itself.
-A malicious guest equipped with such a prediction head can make accurate guesses for labels without accessing other guests’ features.
+If a curious guest inflates its learning rate, the joint model gives preference to its outputs.
+If it can obtain even a small number of labels, it may train its own prediction head.
+The authors show that such a malicious model can make meaningful guesses for labels.
 Luo et al.
@@
 Typical defenses include differential privacy, gradient compression, homomorphic encryption and secret shares for gradients and model features.
-However, these methods usually have a negative effect on model convergence rate and tightness Khan et al.
+However, these methods usually have a negative effect on model convergence rate and tightness (Khan et al., 2023).
+Moreover, implementing such cryptographic algorithms is computationally expensive, and scales poorly.
+3 Related Work
+In this section, we review some closely-related lines of research which have motivated DVFL. A more verbose discussion is available in the Appendix E.
+Fault Tolerant Distributed Optimization
+Like any other distributed system, redundancy enables fault resilience and tolerance in distributed optimization (Liu et al., 2021).
+Liu (2021); Bouhata et al.
+(2022) provide extensive reviews of Byzantine fault tolerant horizontally distributed learning algorithms.
+Most methods hurt model performance and carry significant computational overhead.
+For VFL: For VFL, there is limited work on crash faults.
+However, there is a line of work on the related case of straggling guests.
+That problem definition only extends to the count of failures being bounded as in FedVS (Li et al., 2023), and/or late messages eventually arriving (Xia et al., 2023; Shi et al., 2022; Zhang et al., 2021; Hu et al., 2019).
+A real system could have an arbitrary number and duration of crashes, so these methods do not generalize to crash faults.
+Typically, they do not address faults in the backward pass.
+Paradigms that require periodic synchronization e.g.
+Gradient Assisted Learning (Diao et al., 2022) are similarly hamstrung.
+Asynchronous VFL
+There are some efforts in speeding up training by relaxing synchronization in and splitting effort of gradient calculation.
+But many of these methods expect all inputs during the forward pass (Wang et al., 2024; Shi et al., 2024; Zhang et al., 2021; Gu et al., 2021).
+Chen et al.
+(2020a) present a VFL algorithm which has an asynchronous forward pass, which is a viable candidate for crash tolerance.
+Another line of work (Zhang et al., 2021; Gu et al., 2021) has an implicitly asynchronous forward pass– its aggregation strategy is summation rather than concatenation.
+Federated Transfer Learning
+Federated Transfer Learning (FTL) is a strategy for systems with overlap in both sample space and feature space.
+In Liu et al.
+(2020); Feng & Yu (2020); Feng et al.
+(2022), parties learn a common latent space using local feature extractors and transfer to their tasks.
+For vertically partitioned datasets with small $|\mathcal{S}_{\mathrm{labels}}|$ but large $|\mathcal{S}_{\mathrm{guests}}|$, Sun et al.
+(2023) introduce Few-Shot VFL.
+A few epochs of training on the aligned dataset are followed by clustering-based semi-supervised learning on the unlabelled entities of each client’s dataset.
+Besides the unwanted computational overhead required to generate the labels, this is effectively a label inference attack.
+Greedy and Localized Learning
+Recently, there has been a resurgence of interest in algorithms which train modules of NN layers on local objectives as opposed to BP.
+One reason is inherent locking in BP, described in Jaderberg et al.
 (2017).
-Moreover, implementing such cryptographic algorithms is computationally expensive, and the cost increases with the number of participants.
-3 Related work
-3.1 Fault-Tolerant Distributed Optimization
-The impact of redundancy for fault tolerance in distributed systems is well studied.
-Liu et al.
-(2021) provide a rigorous analysis of redundancy in the case of distributed optimization, showing that systems which are $2f$-redundant are approximately resilient to $f$ Byzantine agents.
-The authors also present two fault-tolerant algorithms for horizontally distributed convex optimization.
-Liu (2021); Bouhata et al.
-(2022) extensively review approaches to fault tolerance in distributed learning.
-Most methods require synchronization by an aggregator and incur computational overhead.
-Further, most methods involve manipulating gradients, which forces a choice between convergence rate and tightness.
-For VFL: Few methods exist for to address fault tolerance in VFL.
-Li et al.
-(2023) introduce FedVS, an algorithm for systems where guest models are polynomial networks Livni et al.
-(2014a; b).
-FedVS uses a form of secret sharing to introduce redundancy in the input data, and can tolerate a fixed number of late or failed guests.
-However, FedVS and many of its predecessors Shi et al.
-(2022); Chen et al.
-(2020a); Hu et al.
-(2019) only address faults in the forward pass.
-3.2 Federated Transfer Learning
-Federated Transfer Learning (FTL) is a strategy for systems where there is some overlap in both sample space and feature space.
-The host owns labels and some features (i.e.
-it is an active participant) for entities in its sample space.
-Guests have limited labels for their task, but have some small overlap in feature space and sample space with the host (“aligned samples”).
-Liu et al.
-(2020) present Secure Transfer Federated Learning (STFL), a framework for a single host and a single guest.
-Both parties learn a common latent space using local feature extractors.
-Encodings of the passive participant’s unaligned samples in this feature space can be used by the passive participant to learn its task model.
-Feng & Yu (2020); Feng et al.
-(2022) extend this idea to multiple parties.
-For the related case of VFL with small $|\mathcal{S}_{\mathrm{guests}}\cap\mathcal{S}_{\mathrm{labels}}|$, Sun et al.
-(2023) introduce few-shot Vertical Federated Learning, where a few epochs of training on the aligned dataset is followed by semi-supervised learning on the non-intersecting entities of each client’s dataset.
-To generate labels, guests cluster the gradients they receive from the aligned samples.
-While this enables remarkable model performance, there is unwanted computational overhead required to generate the labels.
-Moreover, this is a form of label inference attack!
-Not only is this a breach of privacy to some degree, this clustering could be used downstream by a malicious agent to perform feature inference attacks (see Section 2.4.3).
-3.3 Greedy and Localized Learning
-Algorithms which update parameters in hidden layers of ANNs based on local or auxiliary objectives (as opposed to the global label loss) have recently seen renewed interest.
-In the scope of VFL training, they enable some useful properties for ANN training.
-The primary drawbacks of BP in the scope of VFL are the inherent locking synchronization constraints, described in Jaderberg et al.
-(2017).
-In order to update its weights, a layer must wait for the backward signal to reach from the output layer (backward locking).
-In order to begin updating any layer, a full forward pass must first propagate to the end of the network (update locking).
-In order for a layer to evaluate an input to begin with, a forward pass must propagate to it (forward locking).
-These strict locks at the split layer are exactly the failure points for VFL training– forward locking causes training to fail if a guest does not communicate activations, backward locking causes training to fail if the host does not communicate gradients.
-By eschewing the wait to evaluate the global objective, most greedy algorithms are backward unlocked if not update unlocked.
-In order to partially address forward locking for model parallel training Belilovsky et al.
-(2020) propose a replay buffer mechanism.
-Each agent responsible for training a module uses a replay buffer to store activations from the previous module’s agent.
-During a forward pass, the agent reads an activation from the buffer.
-If the forward signal has propagated through the previous module, it would have written the latest activation to the buffer.
-If not, the previous activation is reused.
+With local supervision, a layer need not wait for the global backward signal to update its model (“backward unlocking”).
+Belilovsky et al.
+(2020) propose a replay buffer: Each module of layers stores its inputs in a replay buffer and forward passes its front.
+It need not wait for a fresh forward signal to compute an output (“forward unlocking”).
+Note that forward and backward locking at the split layer are exactly the failure points for VFL.
+Local updates have proved useful for practical HFL (e.g.
+Collins et al.
+(2022); Jeon & Kim (2022); Briggs et al.
+(2020)), where labels are available to guests.
+For VFL, Fu et al.
+(2022b) use cached statistics to construct local gradients.
+As far as we know, DVFL is the first with fully localized model training.
+4 Why Not an Engineering-Level Solution?
+Keen observers will note that the critical failures of the “wait” and “timeout” strategies discussed in §2.4.1 can be partially mitigated with some extra bookkeeping.
+If the guest sends back the sample index (or indices of samples in its batch) and epoch number with each message, the host can uniquely identify which request it belongs to.
+Now timeout may be employed with a strategy to fill missing inputs; late messages may be identified and discarded.
+Wait may be employed by periodically repeating unfulfilled requests; messages from other iterations may be queued.
+So why DVFL?
+Notably, neither strategy handles faults on on backward passes.
+Moreover, wait is a busy-wait.
+Besides being incompatible with deadline-based scheduling and wasting computational and communication bandwidth with repeated unserviced requests, there is no guarantee that training ever continues.
+In fact, this would also be the case for event-driven computation.
+But it is not so for DVFL.
+As for timeout, it is not trivial to choose a good replacement for missing inputs.
+We investigate some obvious options and show that DVFL outperforms them.
 5 Proposed Method
@@
 The first is an unsupervised and decoupled greedy guest-host representation learning phase.
-In this phase, the guest and host models learn feature extractors.
+In this phase, the guest and host models learn feature extractors on their own unsupervised objectives.
 The second is a supervised owner transfer-learning phase where the label owner learns to aggregate encodings from the host space and produce label predictions.
-Within the guest-host phase, participants train their models asynchronously on their own obejctives.
-This allows some flexibility regarding when participants communicate and when they train.
-Since guests do not rely on feedback from hosts to update their models, host or communication faults are irrelevant to the guest models.
+We do not consider faults in owner training.
+They do not affect the guests and hosts which train the bulk of the network.
+Even if the owner were unstable, it only needs to collect hosts’ final encodings for each input once– then, it may train its model offline whenever it is up.
+[^2] This raises the question– why not do this between the guest and the host as it is in FedOnce (Wu et al., 2022)?
+Firstly, the owner’s job is only to predict, so a deep model is not required.
+As such, the dimensions of the embeddings and the computational cost of training its model will typically be small enough to neglect the cost of retrial, unlike a FedOnce/VFL host.
+Moreover, if the aggregator only gets one round of communication as in FedOnce, model performance is hindered.
+We explore this in our experiments and the Appendix.
+Since guests do not rely on feedback from hosts to update their models, host or connection faults during the backward pass are irrelevant to the guest models.
 Hosts store activations from guests as they arrive and use the latest message for their forward pass, so a missing input on the forward pass is no longer critical.
-After hosts and guests complete training, the label owner trains its model.
-For each entity in $\mathcal{S}_{\mathrm{labels}}$, the owner requests guests and hosts to send its encoded and aggregated features.
-The owner uses this as an input for its task model, e.g.
-a classifier.
-We present a DVFL approach to training split neural networks in the follwing subsections.
-The resultant algorithm is formalized in Algorithm 1.
-We also show how DVFL guests might exploit samples outside the joint sample space $\mathcal{S}_{\mathrm{guests}}$ to train better guest and host models.
+The lack of end-to-end feedback combined with the three-tier hierarchy leads to significant robustness.
+The task of a host’s model is only to learn an encoding of features in the joint guest domain, which is not an instance-specific task.
+Therefore, a host may still update its model meaningfully with some out-of-date activations.
+Similarly, guests may use their entire datasets as opposed to only members of the intersection.
+Besides leading to better models, using more data means a single fault does not affect the overall uptime as much.
+On the backward pass, faults only affect the crashed participant.
 5.1 The DVFL Hierarchy
-Guests: The role of a guest in DVFL is identical to a VFL guest’s (see Section 2.1), i.e.
+Guests: The role of a guest in DVFL is identical to a VFL guest’s (see §2.2), i.e.
 guests $g_{i}\in\mathcal{G}$ are the agents $g_{i}$ who own datasets $\mathcal{D}_{i}$ with records of features $\mathcal{F}_{i}$ for the set of entities whose indices are present in $\mathcal{S}_{i}$.
-Each guest has a guest model and an unuspervised objective.
+The guest model’s role is still to encode its dataset’s private features.
+Unlike VFL, they learn to do so on local unsupervised objectives.
 Hosts: Hosts are members of the set $\mathcal{H}$ of agents $h_{i}$ which own aggregating host models, which accept the concatenation of all guest model outputs as input.
@@
   3. 3.
-    Input replay buffer $\mathbf{A}_{i}$. The host $h_{i}$ uses this to store a history of model inputs, if it wishes to reuse them.
-Each host model has the same architecture but its own independent parameters $\theta_{h,i}$. It is possible that $\mathcal{H}\cap\mathcal{G}\neq\phi$.
+    Input replay buffer $\mathbf{A}_{i}$. The host $h_{i}$ uses it to store a history of model inputs if it wishes to reuse them.
+Each host model has the same architecture but its own independent parameters $\theta_{h,i}$.
+Unlike standard VFL, the hosts do not make predictions.
+Their models are only responsible for aggregating guest outputs and learning to effectively map them to a lower dimensional space.
+Note that when there is no such compression, the host can achieve this without any learnable parameters– only a single aggregating layer e.g.
+concatenation is necessary.
 Owner: The owner agent $o$ executes each training and communication round.
-It has access to ground truth labels, and is responsible for training a “final” or “owner” model on concatenated representations from members of $\mathcal{H}$ (i.e.
+It has access to ground truth labels, and is responsible for training a “final” or “owner” model $m_{o}(\text{ }\cdot\text{ };\theta_{o}):\mathbb{R}^{k_{h,1}}\times\dots R^{k_{h,|\mathcal{H}|}}\to\mathbb{R}^{\text{out}}$ on representations from members of $\mathcal{H}$ (i.e.
 it is a transfer learning model).
-The owner has the following:
-  1. 1.
-    Labels $L=\left\{\mathbf{y}_{j}\in\mathbb{R}^{\text{out}}:\text{label }\mathbf{y}_{j}\text{ of }\mathbf{x}_{j},j\in\mathcal{S}_{\mathrm{labels}}\right\}$:
-    • $\mathcal{S}_{\mathrm{labels}}$ is the set of sample indices for which the host has the corresponding labels
-    • We assume that the intersection of all the participants’ indices is not null to allow for training, i.e., $\left|\mathcal{S}_{\mathrm{guests}}\cap\mathcal{S}_{\mathrm{labels}}\right|>0$.
-  2. 2.
-    Final model $m_{o}(\text{ }\cdot\text{ };\theta_{o}):\mathbb{R}^{k_{H}\cdot{|\mathcal{H}|}}\to\mathbb{R}^{\text{out}}$ with parameters $\theta_{o}$.
-The label owner may be a member of $H\cup G$. In any case, we refer to it shorthand as “the owner”. Figure 2 illiustrates a DVFL system with three guests and two hosts.
-Figure 2: A DVFL system with three guests and two hosts.
+The owner has access to the ground truth targets $L=\left\{\mathbf{y}_{j}\in\mathbb{R}^{\text{out}}:\text{label }\mathbf{y}_{j}\text{ of }\mathbf{x}_{j},j\in\mathcal{S}_{\mathrm{labels}}\right\}$.
+Figure 2: A DVFL system with three guests (a) and two hosts (b).
 Guests train their local models on unsupervised objectives and hosts also train their aggregating models on unsupervised objectives.
-After that, the label owner trains a transfer learning model (such as a linear classifier head) on the encodings from the hosts.
+After that, the label owner (c) trains a transfer learning model (such as a linear classifier head) on the encodings from the hosts.
 5.2 Asynchronicity in DVFL
-To avoid a single point of failure, host models should continue training in the absence of inputs from some guests.
-We adapt the use of activation replay buffers of Belilovsky et al.
-(2020) to facilitate training under guest and communication faults.
-We use single-reader single-writer “communication” registers so that hosts can reuse old activations from failed guests.
-A host reads the latest messages in its registers and uses them to fulfill its model inputs.
-If hosts wish to train for more iterations than they receive activations for, they may store their model inputs in replay buffers.
-They may then reuse inputs as needed.
-We discuss these mechanisms in further detail in Appendix D.
+To avoid a single point of failure, host models should continue training without inputs from some guests.
+Single-reader single-writer “communication” registers allow hosts to reuse old activations under guest and communication faults.
+Hosts can store model input history in replay buffers for reuse if they wish to train for more iterations than they receive activations for.
+We discuss these mechanisms in further detail in Appendix F.
 The task of a host’s model is only to learn an encoding of features in the joint guest domain, which is not an instance-specific task.
@@
 5.4 DVFL for Split Training of Neural Networks
-Given the DVFL hierarchy, ANNs are natural choices for participants’ models.
-Unsupervised neural networks are feature extractors Intrator (1992); Becker & Plumbley (1996), which is precisely the role of guest and host models.
-Similarly, the strong task performance of ANNs have been the primary driver of deep learning over the past few years, so they are suitable for owners, too.
-Indeed, just as SplitNN is a popular implementation of VFL, we expect ANNs to be common models for DVFL participants.
-Figure 1 illustrates the relationship between DVFL and VFL in the context of a global ANN model.
-We may say $m_{i}$s are all ANNs with parameters $\theta_{i}$, $m_{h,j}$ are also ANNs whose input is the concatenation of the outputs of all $m_{i}$s, and $m_{o}$ is an ANN whose input is the concatenation of the outputs of all $m_{h,j}$s.
-Algorithm 1 DVFL training for split neural networks
+NNs are natural choices for participants’ models.
+Unsupervised neural networks are feature extractors (Intrator, 1992; Becker & Plumbley, 1996), which is precisely the role of guest and host models.
+Similarly, the strong task performance of NNs have been the primary driver of deep learning over the past few years, so they are suitable for owners, too.
+Indeed, just as SplitNN is a popular implementation of VFL, we expect NNs to be common models for DVFL participants.
+Figure 1 illustrates the relationship between DVFL and VFL in the context of a global NN model.
+We may say $m_{i}$s are all NNs with parameters $\theta_{i}$, $m_{h,j}$ are also NNs whose input is the concatenation of the outputs of all $m_{i}$s, and $m_{o}$ is an NN whose input is the concatenation of the outputs of all $m_{h,j}$s.
+The pseudocode is presented in the Appendix A (Algorithm 1).
+6 Experiments
+Datasets
+As a toy example, we consider a system of $|\mathcal{G}|$ guests which can each see $\frac{1}{|\mathcal{G}|}^{th}$ of an handwritten digit from the MNIST dataset (Deng, 2012).
+For our experiments, we set $|\mathcal{G}|=4$ i.e.
+$g_{1}$’s dataset contains the top $28\times 7$ pixel patch of an MNIST image, $g_{2}$’s dataset contains the second , $g_{3}$’s contains the third, and $g_{4}$’s contains the bottom $28\times 7$ patch (see Fig.
+3(a)).
+We also experiment on some benchmark VFL datasets in the Appendix.
+New Baselines and Loss Functions
+DVFL is an alternate strategy to VFL. Naturally we compare DVFL for NNs to SplitNN, a direct application of VFL to NNs. Naive VFL implicitly employs the “wait” protocol discussed in §2.4.1.
+We also try the modified timeout strategy from §4 by imputing missing inputs with zeros (zeros), stale activations (buffer), and skipping that round altogether (skip).
+For SplitNN hosts and DVFL owners, we use cross-entropy loss.
+For DVFL hosts and guests, We use a learnable decoder head and MSE reconstruction as an unsupervised objective.
+FedOnce Baseline
+We also provide a baseline with FedOnce (Wu et al., 2022).
+FedOnce is similar to DVFL in the sense that guests train their own models on unsupervised objectives, and similar to VFL in the sense that there is a centralized aggregating host that owns labels.
+FedOnce only requires one round of communication (the guest-encoded dataset after all guests have fully trained).
+It can be argued that each guest can retry until the host has heard from all of them (since it would not be holding up training of any future iterations).
+If that is implemented, faults are not a worry.
+Regardless, we also try an alternate fault-handling strategy inspired by VFL-zeros for compatibility with deadline-based scheduling.
+Hyperparameters, model details and ablation study are presented in the Appendix.
+Asynchronous Baselines
+We also compare DVFL to the two discussed viable asynchronous VFL algorithms: VAFL (Chen et al., 2020a) and VFB2 (Zhang et al., 2021).
+For VAFL, we eschew random batch sampling for iteration (in the same order by all guests).
+This locality gives us a skyline of VAFL’s performance– non-active guests’ “stale” activations will be recent.
+For VFB2, we simulate without the two distinct communication graphs to minimize the scope for faults.
+Further, since we experiment with only one label-owning party, the “active parties” class only contains a single host ($m=k=1$ by the convention of the original paper).
+Also, since the aggregation strategy is summation, we modify the first layer of the host model’s input dimension to equal a single host’s output dimension.
+FedVS Baseline and More Datasets
+We also compare the model performance of SplitNN and DVFL to results cited in the work FedVS (Li et al., 2023).
+Since FedVS is incompatible with our fault model, we only compare performance under perfect conditions.
+We use the same datasets (Yeh & hui Lien (2009); Markelle Kelly ; Li et al.
+(2022); Sakar et al.
+(2019)) as used in the experiments for FedVS.
+FedVS restricts its models to use polynomial networks (Livni et al., 2014a) i.e.
+equally-sized linear layers with polynomial activation functions.
+The models we use for SplitNN and DVFL are smaller– each layer is half as wide as the next, and we use Leaky ReLUs.
+6.1 Fault Simulation
+Six hyperparameters control the fault model described in §2.4.1.
+$R^{(d)}_{i}$ is the fault rate for a live $i\in\{\mathrm{connection},\mathrm{guest},\mathrm{host}\}$.
+$R^{(u)}_{i}$ is the chance a dead $i$ rejoins the network before the next request.
+We discuss how exactly the faults are simulated in the Appendix.
+We assume no faults during label supervision or inference.
+6.2 Experimental Setup
+Table 1: We compare DVFL to standard VFL.
+After filtering outliers by IQR, we report the mean and 2x the standard deviation of MNIST test performance when trained with various entities $i\in\{\mathrm{connection,guest,host}\}$ susceptible to fail at a rate of $R^{(d)}_{i}=0.3$.
+A failed resource has $R^{(u)}_{i}=\mathbb{P}_{\mathrm{rejoin}}\in\{1,0.5,0.1\}$ to be available next time it is requested.
+SplitNN, the baseline VFL algorithm, cannot train at all when there are any faults.
+On the other hand, our DVFL implementation degrades gracefully with faults.
+Using an explicit fault handling strategy within VFL does allow it to train.
+But the degradation with faults is sharper than DVFL and the variance in model performance is higher.
+Under fault-free circumstances, SplitNN slightly outperforms our DVFL implementation.
+With minor adjustments, SplitNN can tolerate some faults, but our approach is more resilient and stable.
+Strategy Test Accuracy (%) with 0.3 Loss Rate
+No Faults Connection Loss Guest Loss Host Loss
+$\mathbb{P}_{\mathrm{rejoin}}=$ N/A $\mathbb{P}_{\mathrm{rejoin}}=$ 1 $\mathbb{P}_{\mathrm{rejoin}}=$ 0.5 $\mathbb{P}_{\mathrm{rejoin}}=$ 0.1 $\mathbb{P}_{\mathrm{rejoin}}=$ 1 $\mathbb{P}_{\mathrm{rejoin}}=$ 0.5 $\mathbb{P}_{\mathrm{rejoin}}=$ 0.1 $\mathbb{P}_{\mathrm{rejoin}}=$ 1 $\mathbb{P}_{\mathrm{rejoin}}=$ 0.5 $\mathbb{P}_{\mathrm{rejoin}}=$ 0.1
+DVFL-NN 97.80$\pm$0.12 97.78$\pm$0.11 97.78$\pm$0.14 97.72$\pm$0.17 97.77$\pm$0.11 97.74$\pm$0.16 97.58$\pm$0.20 97.81$\pm$0.12 97.79$\pm$0.14 97.60$\pm$0.33
+VFL (SplitNN) 97.85$\pm$0.17 N/A N/A N/A N/A N/A N/A N/A N/A N/A
+VFL (SplitNN-skip) 97.86$\pm$0.16 97.50$\pm$0.54 96.60$\pm$0.51 77.21$\pm$5.34 97.44$\pm$0.57 96.65$\pm$0.45 80.22$\pm$4.31 97.65$\pm$0.41 97.59$\pm$0.50 97.29$\pm$0.89
+VFL (SplitNN-zeros) 97.84$\pm$0.15 97.64$\pm$0.38 97.53$\pm$0.40 97.44$\pm$0.36 97.68$\pm$0.40 97.59$\pm$0.40 96.95$\pm$0.50 97.68$\pm$0.30 97.65$\pm$0.37 97.31$\pm$0.93
+VFL (SplitNN-buffer) 97.85$\pm$0.17 97.66$\pm$0.38 96.85$\pm$0.74 93.45±$\pm 2.10$ 97.59$\pm$0.58 96.76$\pm$0.70 91.78$\pm$1.70 97.81$\pm$0.32 97.71$\pm$0.40 97.69$\pm$0.51
+VFL (FedOnce-zeros) 97.76$\pm$0.20 96.89$\pm$0.29 96.41$\pm$0.30 92.80$\pm$0.75 97.47$\pm$0.37 97.45$\pm$0.31 97.50$\pm$0.31 97.28$\pm$0.35 97.34$\pm$0.37 96.93$\pm$0.51
+VAFL (skyline) 97.57$\pm$0.30 97.17$\pm$0.46 96.88$\pm$0.36 96.46$\pm$0.44 97.18$\pm$0.42 96.99$\pm$0.44 96.45$\pm$0.51 97.17$\pm$0.48 96.89$\pm$0.36 96.42$\pm$0.52
+VFB2 (skyline) 97.72$\pm$0.16 97.43$\pm$0.38 97.33$\pm$0.41 96.09$\pm$0.47 97.36$\pm$0.45 97.24$\pm$0.41 96.29$\pm$0.42 97.43$\pm$0.49 97.20$\pm$0.61 96.79$\pm$0.90
+Performance under Faults
+In order to show that it is possible to use DVFL even with crash faults, we train DVFL with $|\mathcal{H}|=4$ hosts with a fault rate of 0.3 and a rejoin rate of 0.1.
+For a given experiment, the failure mode (connection, guest, or host) is fixed.
+$R^{(u)}=1$ simulates single-message loss, since the next time a message is sent or received, the participant is sure to have rejoined.
+$R^{(u)}=0.5$ simulates short bursts of loss.
+$R^{(u)}=0.1$ simulates appreciable downtime.
+Effect of Redundancy
+In order to show that redundancy of hosts improves model performance and fault tolerance, we simulate DVFL training with $|\mathcal{H}|$ = 1, 2, 3, and 4 hosts.
+We measure test accuracy given connection fault rate $\in\{0,0.3,0.6\}$ during training.
+Learning from a Limited Intersection
+We first fix a small labelled and aligned intersection size.
+The remaining sample space is shuffled and distributed evenly across all four guests.
+For example, if we wish to use $1024$ labeled samples, we shuffle samples indexed $1024-59999$ and split the $58976$ samples evenly, such that each guest gets $14744$ unique MNIST patches + $1024$ intersecting MNIST patches = $15768$ total samples.
+We try this experiment with $128$, $256$, $512$ and $1024$ labeled samples.
+We compare with SplitNN trained on only the labeled samples.
+7 Results and Discussion
+(a) For our experiments, the input received by each guest for a given sample, as well as the reconstruction its autoencoder model produces.
+[…]
+(b) Model performance improves with redundancy, especially when there is a nonzero crash rate.
+7.1 Graceful Degradation with Faults
+(Table 8) shows that model performance degrades with faults, but not catastrophically (unlike VFL with SplitNN).
+In ideal conditions, DVFL performance is slightly poorer than SplitN’s.
+This is not unexpected as we will discuss in §9 Observe that the degradation with faults in DVFL is much more graceful and stable than the SplitNN variants.
+In the case of SplitNN-based algorithms, a connection and guest fault are functionally identical.
+In the case of DVFL, there are some performance savings under connection faults– the corresponding guest and host can still meaningfully learn even in the absence of connection.
+Further, redundant hosts may continue training as normal.
+Tolerating some faults pays off
+The artifical input induces a Dropout effect (Srivastava et al., 2014).
+This reduces bias but competes with information loss in the model.
+DVFL models, without cross-device feedback, see smaller information loss compared to SplitNN-based methods (which do not respond to faults as well, but still exhibit somewhat graceful test accuracy degradation).
+The simplest SplitNN-based method, Skip, leads to the worst performance since information from non-faulty guests is also lost.
+Zeros works as if failed neurons do not exist at all in simple linear and convolutional layers.
+This would fail with more complex architectures and algorithms e.g.
+skip connections and constrained solvers.
+Buffer empirically cannot exploit locality in activations and performs poorly.
+Reconstructing missing inputs based on statistics, as Fu et al.
+(2022b) do for gradients, may be the best solution but would suffer from bias and computational infeasibility.
+Comparison to asynchronous VFL
+Asynchronous algorithms suffer from higher variance, perhaps since the information loss from a single failure do not have the safety net of information from successful interactions.
+Further, these skylines perform significantly worse than VFL and DVFL even when there are no faults.
+In fact, we argue that VFL-zeros provides a skyline for performance from any asynchronous algorithm (at least when the model has no skip connections or optimization constraints).
+As discussed in the previous paragraph, VFL-zeros acts as if failed neurons did not exist at all.
+This is as if an oracle told an asynchronous host which guests/connections will be live, and the host calculated losses based on multiple guests’ embeddings at once hence leading to lower-variance loss estimates.
+VFB2 and other algorithms using summation-based aggregation also suffer from constraints on the model architecture.
+The output of each guest must have the same dimension.
+Further, the host input dimension is smaller than what it would be if it were to aggregate by concatenation, so there are fewer degrees of freedom for the model.
+Comparison to FedVS and results on other datasets
+FedVS (Li et al., 2023) provides both gradient privacy (via cryptographic multi-party communication) and straggler resistance.
+The information loss thanks to quantization and random masking in FedVS outweighs that from localized training in DVFL.
+As a result, performance lags even in perfect conditions (Table 2).
+We also see that DVFL performance tracks SplitNN’s closely on all datasets, including tabular ones which are typically “hard” for neural networks.
+Table 2: On a variety of typical vertically partitioned datasets, we see that DVFL performs comparably to SplitNN even under perfect conditions.
+Even on tabular datasets harder (especially for NNs) than MNIST such as Credit Card and Parkinson’s, DVFL model performance is comparable to SplitNN.
+Despite using a smaller model, both outperform the results cited in the FedVS paper (Li et al., 2023).
+Strategy Test Accuracy (%) Desiredata
+Dataset Privacy Fault Tolerance
+Credit Card Parkinson’s CalTech-7 Handwritten
+DVFL 82.66 85.76 99.00 98.25 No Gradients Crash
+VFL (SplitNN) 84.11 86.64 99.66 98.25 None None
+FedVS (Li et al., 2023) [^3] Results extracted from graphs. Raw figures were unavailable and the authors did not respond to our inquiries. 81.99 81.83 95.98 98.25 Quantization and Noise Straggler
+005555101010101515151520202020252525259797979797.297.297.297.297.497.497.497.497.697.697.697.697.897.897.897.8Guest training epochs per communication epochMNIST Test Accuracy (%)
+Figure 3: Model performance degrades with an increase in communication period, i.e. more communication is correlated with better performance.
+Comparison to FedOnce
+FedOnce (Wu et al., 2022) aims to reduce the volume of communication at the cost of model performance.
+Equivalently, Belilovsky et al.
+(2020); Siddiqui et al.
+(2023) demonstrate that updating all layers on local objectives on each forward pass performs better than fully training and freezing a layer before moving on to the next.
+Even if we consider this completely fault tolerant, the overall performance loss from reduced communication is enough that DVFL performs better or comparably even in our experiments with the high fault rates.
+If communication size is a concern, DVFL provides fine-grained control over that.
+This is discussed further in §8.1.
+7.2 Redundancy Improves Performance, Asynchronicity Affords Efficiency
+We observe (Figure 3(b)) that test accuracy improves with the number of hosts with all other system parameters fixed.
+This is the case even when there are no faults.
+This is not surprising, as the owner’s aggregation of host results can be seen as a form of bagging (Breiman, 1996).
+As a corollary, the degradation of performance with connection faults is more graceful with increasing redundancy.
+7.3 Out-of-Intersection Training and Privacy
+Although VFL with SplitNN marginally outperforms our DVFL NN in the experiments with perfect conditions, DVFL significantly outperforms VFL on a limited labeled intersection.
+DVFL’s advantage is greater with fewer labels (Table 3).
+DVFL is performant while also completely eliminating the possibility of gradient-based inference attacks.
+As reviewed in 3, complex mechanisms can be developed on top of SplitNN to provide privacy guarantees within some bounds.
+DVFL achieves complete privacy against gradient-based attacks without any extra computation and minimal performance loss.
+Table 3: We exploit data outside the sample intersection to train DVFL’s guests and hosts.
+For the owner model, we use labeled entities common to all datasets.
+This performs better than VFL, which can only use the common samples.
+Strategy Test Accuracy (%)
+# of Labeled Samples
+128 256 512 1024
+DVFL 76.45 82.51 85.18 88.90
+VFL (SplitNN) 65.07 77.51 83.09 87.26
+8 Engineering Flexibility with DVFL
+8.1 Control over Communication Volume
+Since guests and hosts both train on their own objectives, it is not necessary that they synchronize on every iteration.
+This means communication cost can be controlled by only transferring some embeddings, at the cost of performance.
+We experiment with various communication frequencies and present the results in Table 3.
+These experiments are discussed in further detail in the Appendix F.1.
+8.2 Control over Computational Cost
+Ostensibly, DVFL is slightly more computationally expensive than VFL since both guests and hosts must generate their own update signals in the absence of supervisory labels.
+In our experiments with autoencoders, this manifests as training the decoder head.
+However, DVFL allows the flexibility of allowing each participant to train for a different number of iterations.
+This is unlike VFL where the guest must compute a forward pass and then update every time the host makes a request.
+As such, the number of iterations for each participant can be tuned to balance model performance and computational cost.
+Our main experiments with DVFL only use 20 epochs of guest training while VFL hosts, DVFL hosts and DVFL owners use 60 epochs.
+We measure the wall-clock time of each process in the Appendix F.3 and see that the guest processes are actually alive for a shorter duration in DVFL in comparison to VFL ($\approx 55.88$s vs $\approx 80.22$s under the same operating conditions).
+Indeed, the same way that the meaningful learning is possible even if the guest sends a limited number of activations, the host may make do with storing a limited number of activations.
+In the Appendix F.2 we discuss how a memory constrained host may still meaningfully learn while only storing the last few activations rather than the entire history.
+9 Limitations
+Our models are shallow.
+Wang et al.
+(2021) show that the performance improves with the length of backpropagation in blockwise-trained networks, and suggest that it is due to increasing mutual information between the shallower and deeper layers.
+Further, given a model, the loss does not guide the model to extract strictly task-useful features.
+These issues typically also lead to scalability issues to harder datasets, although we observe in our experiments on tabular datasets that the difference is not that large.
+However, we expect that the information loss from the absence of direct connections between guests’ neurons is greater than that due to the loss of task-relevant information.
+10 Broader Impact
+We believe DVFL will enable large scale and democratic participation in joint learning tasks where VFL has been unable to provide solutions.
+e.g.
+street & private residential cameras predicting traffic incidents at a busy intersection.
+Participants might not all be able to see each training incident i.e.
+small $\mathcal{S}_{\text{guest}}\cap S_{\text{labels}}$, or might be concerned about privacy, or suffer from intermittent connection.
+DVFL is also more practical for existing cross-silo applications.
+Consider the typical example of various specialized healthcare providers predicting, say, cancer.
+Not only can the hospitals train asynchronously and not worry about faults, they need not worry about compromised hospitals launching gradient inference attacks.
+Further, they may now admit more hospitals to their system even if they have few patients in common.
+We see DVFL as a step towards usable and ubiquitous machine learning.
+11 Conclusion and Future Work
+We present Decoupled Vertical Federated Learning, a strategy for joint learning on vertically partitioned data.
+Instead of adding complexity to cope with the data leakage and synchronization associated with BP, we eschew feedback altogether.
+Combined with the ability to reuse feedforward signals, this allows for fault tolerance in training.
+All-in-all, we address three impracticalities with VFL training: single points of failure, training data being limited to the intersection, and gradient-based privacy attacks: Training may continue even when there are faults, or there are a limited number of samples common to all guests.
+Malicious or compromised guests can no longer use gradient feedback to figure out possibly sensitive information.
+Existing techniques to enforce privacy guarantees either hamper model convergence, are expensive to implement or both.
+In turn, DVFL introduces new flexibility and practicality to systems intended for learning from vertically partitioned data– the volume of communication, memory usage and overall complexity can be controlled by a designer.
+Our results suggest DVFL’s performance to be comparable to SplitNN on various ML tasks.
+Unlike SplitNN, DVFL degrades gracefully with manykinds of loss faults.
+Even with explicit fault handling strategies, SplitNN’s performance drops off more sharply than DVFL’s.
+In general, DVFL’s asynchronicity provides a great deal of engineering flexibility.
+Some of it is discussed in the Appendix.
+The next step for DVFL is to generalize it for more tasks and general architectures.
+VFL is readily compatible with a variety of machine learning models (including ensemble models).
+With a judicious choice of training objective for guest models, DVFL may also be extended past NNs.
+Appendix A DVFL for Split Neural Network Training
+Algorithm 1 DVFL training for split neural networks (DVFL-NN)
 Input: Guest datasets $\mathcal{D}_{i}$, optimizers $\text{optim}_{g_{i}}$ $\forall i\in$$\{o\}\cup$$\{g_{1},\dots,g_{|\mathcal{G}|}\}\cup\{h_{1},\dots,h_{|\mathcal{H}|}\}$, communication schedule
@@
 8:     Compute an unsupervised loss $\mathcal{L}_{h,i}\left(\hat{\mathbf{b}}_{t,i}\right)$
-9:     $\theta_{h,i}\leftarrow\text{optim}_{h_{i}}(\theta_{h,i},\nabla_{\theta_{h,i}}\mathcal{L}_{h,i})${Update host model}
+9:     $\theta_{h,i}\leftarrow\text{optim}_{h_{i}}(\theta_{h_{i}},\nabla_{\theta_{h_{i}}}\mathcal{L}_{h_{i}})$Update host model}
 10:  end for
@@
 21:  until convergence or a fixed number of iterations
-5 Experiments
-As a toy example, we consider a system of $|\mathcal{G}|$ guests which can each see $\frac{1}{|\mathcal{G}|}^{th}$ of an handwritten digit from the MNIST dataset.
-For our experiments, we set $|\mathcal{G}|=4$ i.e.
-$g_{1}$’s dataset contains the top $28\times 7$ pixel patch of an MNIST image, $g_{2}$’s dataset contains the second , $g_{3}$’s contains the third, and $g_{4}$’s contains the bottom $28\times 7$ patch (see Fig.
-3).
-Figure 3: For our experiments, the input received by each guest for a given sample, as well as the reconstruction its autoencoder model produces.
-We use a learnable decoder head and MSE reconstruction as an unsupervised objective.
-Baseline
-DVFL is an alternate strategy to VFL.
-Naturally we compare DVFL for ANNs to SplitNN, a direct application of VFL to ANNs.
-Naive VFL implicitly employs the “wait” protocol discussed in Sec 2.4.1.
-There are some simple additions to SplitNN that would allow a more meaningful comparison.
-Firstly, if the host sets a deadline, we may continue training despite missing inputs by skipping (we call this SplitNN-skip) or by supplying replacement values.
-For replacement values, we try a zero tensor (SplitNN-zeros).
-Hosts may ask guests to send the identities of the entity to which each of their messages belong.
-This way, a late message cannot be believed to be intended for the next iteration.
-Results on more baselines and datasets are presented in Appendix A.
-The choice of baseline is discussed in Appendix C.
-5.1 Fault Simulation
-Six hyperparameters control the fault model described in Section 2.4.1.
-$R^{(d)}_{i}$ is the fault rate for a live $i\in\{\mathrm{connection},\mathrm{guest},\mathrm{host}\}$.
-$R^{(u)}_{i}$ is the chance a dead $i$ rejoins the network before the next request.
-We discuss how exactly the faults are simulated in Appendix B.2.
-We assume no faults during label supervision or inference.
-5.2 Experimental Setup
-Table 1: We compare DVFL to standard VFL.
-Here, we report the median and twice the standard deviation of MNIST test performance when trained with various entities $i\in\{\mathrm{connection,guest,host}\}$ susceptible to fail at a rate of $R^{(d)}_{i}=0.3$.
-A failed resource has $R^{(u)}_{i}=\mathbb{P}_{\mathrm{rejoin}}\in\{1,0.5,0.1\}$ to be available next time it is requested.
-SplitNN, the baseline VFL algorithm, cannot train at all when there are any faults.
-On the other hand, our DVFL implementation degrades gracefully with faults.
-Using an explicit fault handling strategy within VFL does allow it to train.
-But the degradation with faults is sharper than DVFL and the variance in model performance is higher.
-Under fault-free circumstances, SplitNN slightly outperforms our DVFL implementation.
-With minor adjustments, SplitNN can tolerate some faults, but our approach is more resilient and stable.
-Strategy Test Accuracy (%) with 0.3 Loss Rate
-No Faults Communication Loss Guest Loss Host Loss
-$\mathbb{P}_{\mathrm{rejoin}}=$ N/A $\mathbb{P}_{\mathrm{rejoin}}=$ 1 $\mathbb{P}_{\mathrm{rejoin}}=$ 0.5 $\mathbb{P}_{\mathrm{rejoin}}=$ 0.1 $\mathbb{P}_{\mathrm{rejoin}}=$ 1 $\mathbb{P}_{\mathrm{rejoin}}=$ 0.5 $\mathbb{P}_{\mathrm{rejoin}}=$ 0.1 $\mathbb{P}_{\mathrm{rejoin}}=$ 1 $\mathbb{P}_{\mathrm{rejoin}}=$ 0.5 $\mathbb{P}_{\mathrm{rejoin}}=$ 0.1
-DVFL-NN 97.30$\pm$0.08 97.24$\pm$0.10 97.19$\pm$0.14 97.03$\pm$0.17 97.11$\pm$0.11 97.08$\pm$0.16 96.87$\pm$0.20 97.27$\pm$0.12 97.24$\pm$0.14 97.02$\pm$0.33
-VFL (SplitNN) 97.80$\pm$0.17 N/A N/A N/A N/A N/A N/A N/A N/A N/A
-VFL (SplitNN-skip) 97.84$\pm$0.16 97.50$\pm$0.54 96.60$\pm$0.51 77.21$\pm$5.34 97.44$\pm$0.57 96.65$\pm$0.45 80.22$\pm$4.31 97.65$\pm$0.41 97.59$\pm$0.50 97.00$\pm$0.89
-VFL (SplitNN-zeros) 97.79$\pm$0.15 97.64$\pm$0.38 97.53$\pm$0.40 96.89$\pm$0.36 97.68$\pm$0.40 97.59$\pm$0.40 96.45$\pm$0.50 97.68$\pm$0.30 97.65$\pm$0.37 97.01$\pm$0.93
-Performance under Faults
-In order to show that it is possible to use DVFL even with crash faults, we train DVFL with $|\mathcal{H}|=4$ hosts under $R^{(d)}_{i}=0.3,R^{(d)}_{j}=0$ where $i\in\{\mathrm{connection},\mathrm{guest},\mathrm{host},\phi\},j\neq i$.
-Given a $R^{(d)}_{i}\neq 0$, we try $R^{(d)}_{i}\in\{0.1,0.5,1.0\}$.
-Effect of Redundancy
-In order to show that redundancy in hosts improves model performance and fault tolerance, we simulate DVFL training with $|\mathcal{H}|$ = 1, 2, 3, and 4 hosts.
-We measure test accuracy given $R_{\mathrm{connection}}\in\{0,0.3,0.6\}$ during training.
-Learning from a Limited Intersection
-We first fix a small intersection size, where all members of the intersection are labeled.
-We generate the remaining dataset for each device by shuffling the out-of-intersection samples and taking windows such that the remaining sample space is distributed evenly across all four guests.
-For example, if we wish to use $1024$ labeled samples, we shuffle samples indexed $1024-59999$ and split the $58976$ samples evenly, such that each guest gets $14744$ unique MNIST patches + $1024$ intersecting MNIST patches = $15768$ total samples.
-We try this experiment with $128$, $256$, $512$ and $1024$ labeled samples.
-We compare with SplitNN trained on only the labeled samples.
-6 Results and Discussion
-[…]
-Figure 4: Model performance improves with redundancy, especially when there is a nonzero crash rate.
-6.1 Graceful Degradation with Faults
-We observe (Table 1) that model performance degrades with faults, but not catastrophically (unlike VFL with SplitNN).
-In ideal conditions, DVFL performance is slightly poorer than SplitNN.
-This is not unexpected, since our models are shallow.
-Wang et al.
-(2021) show performance improves with the length of backpropagation in blockwise-trained networks, and suggest it is due to increasing mutual information between shallower and deeper layers.
-Further, given a model, the reconstruction loss guides the feature extractor to features that will be useful for reconstruction– not necessarily good for classification.
-It is expected that the gap between DVFL and SplitNN will reduce with more general loss functions van den Oord et al.
-(2019); Chen et al.
-(2020b); Caron et al.
-(2020); Zbontar et al.
-(2021) and with self-supervised classification objectives.
-Further, observe that the degradation with faults in DVFL is much more graceful than our SplitNN variants.
-Not only does the SplitNN-wait/skip/stale performance drop more sharply, the variance increases more than it does with DVFL.
-Guest faults are the worst
-In the case of DVFL, the most severe type of fault is the guest loss.
-For example, when $R^{(d)}_{\mathrm{guest}}=0.3,R^{(u)}_{\mathrm{guest}}=0.1$ the accuracy drops from 97.30% to 96.87% while with communication and host loss it is 97.03 and 97.02% respectively.
-The performance of the combined model depends on unique outputs from guests– there is no redundancy.
-With a host or communication fault, guests are still able to update their own models (and there might be redundant hosts, too).
-In the case of SplitNN-skip and SplitNN-host, a communication fault is just as bad– a guest cannot send forward passes to (or receive backward passes from) the single host whether it or the connection is down.
-A host fault is equivalent to skipping that round for training.
-$R^{(u)}=1$ simulates single-message loss, since the next time a message is sent or received, the participant is sure to have rejoined.
-$R^{(u)}=0.5$ simulates short bursts of loss.
-$R^{(u)}=0.1$ simulates appreciable downtime.
-Tolerating some faults pays off
-Induced by the artificial input, the Dropout effect Srivastava et al.
-(2014) leads to a reduction in bias.
-This competes against the loss of information in the model.
-In the case of DVFL, models do not accept feedback, so the scope for information loss is smaller.
-In the case of SplitNN with missing input imputation, models do not respond to faults as well as DVFL.
-Still, the test accuracy degrades somwewhat gracefully for this reason.
-Other failure modes
-In this study, we only examine loss faults.
-Other typical failure modes for distributed systems include arbitrary delays, incorrect order, and Byzantine faults.
-For the forward pass, the first two might be mitigated within VFL itself.
-Since hosts request encoded features from a particular sample or batch of samples, they may ask guests to send back the identity of those samples along with the features (as we assumed in the variants of SplitNN.
-While the privacy is dubious, there is precedent in VFL: see Sun et al.
-(2021); Yang et al.
-(2023); Jiang et al.
-(2022).
-But, in order to correctly serialize, the host must maintain a cache of unused messages.
-For the backwad pass, if a guest has sent two outputs before receiving a single gradient, it has no way of knowing which output it corresponds to.
-DVFL is immune to these issues, since guests and hosts train independently to each other.
-6.2 Redundancy Improves Performance
-We observe (Figure 4) that test accuracy improves with the number of hosts with all other system parameters fixed.
-This is not surprising, as the owner’s aggregation of host results can be seen as a form of bagging Breiman (1996).
-As a corollary, the degradation of performance with communication faults is more graceful with increasing redundancy.
-The role of the DVFL host
-Observe that host faults do not hurt DVFL performance very much.
-DVFL hosts’ primary function is aggregation.
-A DVFL host can achieve this without any learnable parameters, by simply concatenating inputs and passing them onto the owner when requested.
-The other role a DVFL host might play is to project the concatenated host output to a smaller dimension.
-What the host model must learn is to preserve useful features in the new dimension.
-Table 2: We exploit data outside the sample intersection to train DVFL’s guests and hosts.
-For the owner model, we use labeled entities common to all datasets.
-This performs better than VFL, which can only use the common samples.
-Strategy Test Accuracy (%)
-# of Labeled Samples
-128 256 512 1024
-DVFL 76.45 82.51 85.18 88.90
-VFL (SplitNN) 65.07 77.51 83.09 87.26
-6.3 Out-of-Intersection Training
-Although VFL with SplitNN outperforms our DVFL ANN algorithm in our experiments with perfect conditions, DVFL significantly outperforms VFL on a limited labeled intersection.
-DVFL’s advantage is greater with fewer labels (Table 2).
-6.4 Security & Privacy
-DVFL achieves these results while also completely eliminating the possibility of gradient-based inference attacks.
-As reviewed in 3, complex mechanisms can be developed on top of SplitNN to provide privacy guarantees within some bounds.
-DVFL achieves complete gradient privacy without any overhead or model convergence issues.
-7 Conclusion and Future Work
-We present Decoupled Vertical Federated Learning, a strategy for joint learning on vertically partitioned data.
-Instead of adding complexity to cope with the data leakage and synchronization associated with BP, we eschew feedback altogether.
-Combined with the ability to reuse feedforward signals, this allows for fault tolerance and privacy in training.
-Training may continue even when there are faults, or there are a limited number of samples common to all guests.
-DVFL also provides a high degree of privacy.
-Malicious or compromised guests can no longer use gradient feedback to figure out possibly sensitive information.
-Existing techniques to enforce privacy guarantees either hamper model convergence, are expensive to implement or both.
-Our results suggest its performance to be comparable to SplitNN on various machine learning tasks.
-Unlike SplitNN, DVFL degrades gracefully with various kinds of loss faults.
-Even with explicit fault handling strategies, SplitNN’s performance drops off more sharply than DVFL’s.
-In general, DVFL’s asynchronicity provides a great deal of engineering flexibility.
-Some of it is discussed in the Appendix.
-The next step for DVFL is to generalize it for more tasks and general architectures.
-VFL is readily compatible with a variety of machine learning models (including ensemble models).
-With a judicious choice of training objective for guest models, DVFL may also be extended past ANNs.
-Another possible direction for DVFL is the introduction of more elaborate fault handling strategies.
-In this work, the aggregating “host” uses the the latest activation received from a “guest” client.
-This allows for a simple implementation using single-reader single-writer registers.
-We observe that this leads to meaningful representation learning– empirically, we only observe a 0.26% drop in accuracy on MNIST even when there is a 30% crash fault rate.
-In a bid to close this gap further, hosts might build a model of the distribution of its inputs.
-Hosts could then sample missing inputs from its posterior conditioned on successful messages.
-A natural first step would be to use a running mean as an estimator.
-Appendix A Results on More Datasets, Comparison to FedVS
+Appendix B Results on More Datasets, Comparison to FedVS
 In order to bolster the comparison to SplitNN, we evaluate the performance of DVFL on some benchmark vertically partitioned datasets.
-They are Credit Card Yeh & hui Lien (2009), Parkinson’s Sakar et al.
-(2019), CalTech-7 and Handwritten Markelle Kelly .
-Note that CalTech-7 is a subset of 7 classes CalTech 101 Li et al.
-(2022): Face, Motorbikes, Dolla-Bill, Garfield, Snoopy, Stop-Sign and Windsor-Chair.
+They are Credit Card (Yeh & hui Lien, 2009), Parkinson’s (Sakar et al., 2019), CalTech-7 and Handwritten (Markelle Kelly, ).
+Note that CalTech-7 is a subset of 7 classes CalTech 101 (Li et al., 2022): Face, Motorbikes, Dolla-Bill, Garfield, Snoopy, Stop-Sign and Windsor-Chair.
 Both Handwritten and CalTech have 6 distinct “views”– features of distinct domains, e.g.
 spatial, frequency (Fourier), Gabor filter coefficients, etc.
-FedVS Li et al.
-(2023) is a state-of-the-art VFL algorithm which aims to tackle both privacy and fault tolerance.
+FedVS (Li et al., 2023) is a state-of-the-art VFL algorithm that aims to tackle both privacy and fault tolerance.
 There is a round of secret shares so that each guest holds an encrypted version of a horizontal section of other guests’ data and model parameters.
-The shares are arranged in such a manner that if $\leq L$ (a fixed system parameter) guests don’t send their outputs at the forward pass, there is enough redundant information in the other guests’ forward passes to account for it.
+The shares are arranged such that if $\leq L$ (a fixed system parameter) guests don’t send their outputs at the forward pass, there is enough redundant information in the other guests’ forward passes to account for it.
 In order to maintain privacy despite this sharing of data, guests quantize their models and data and apply randomly mask.
@@
 But, this obscurity is at the cost of model performance.
-Further, this algorithm is contingent on the guests using polynomial networks Livni et al.
-(2014a; b).
+Further, this algorithm is contingent on the guests using polynomial networks (Livni et al., 2014a; b).
 In the following experiment, we compare our DVFL algorithm, SplitNN and FedVS.
@@
 We also take the liberty of using non-polynomial activations (Leaky ReLU).
-Model details and hyperparameters are in Appendix B.
+Model details and hyperparameters are in the Appendix.
 $|\mathcal{H}|=1$ for these experiments.
 Table 4: On a variety of typical vertically partitioned datasets, we see that DVFL performs comparably to SplitNN even under perfect conditions.
-Even on tabular datasets harder (especially for ANNs) than MNIST such as Credit Card and Parkinson’s, DVFL model performance is comparable to SplitNN.
-Despite using a smaller model, both outperform the results cited in the FedVS paper Li et al.
-(2023).
+Even on tabular datasets harder (especially for NNs) than MNIST such as Credit Card and Parkinson’s, DVFL model performance is comparable to SplitNN.
+Despite using a smaller model, both outperform the results cited in the FedVS paper (Li et al., 2023).
+DVFL achieves this while tolerating a broader class of faults and also providing complete immunity to gradient-based attacks.
 Strategy Test Accuracy (%)
@@
 VFL (SplitNN) 84.11 86.64 99.66 98.25
-FedVS Li et al. (2023) [^2] Results extracted from graphs. Raw figures were unavailable and the authors did not respond to our inquiries. 81.99 81.83 95.98 98.25
-The feature splits and train-test splits for each dataset are identical to the FedVS paper Li et al. (2023). Further discussion about these results is available in Appendix C.
-Appendix B Models, Hyperparameters and Implementation Details
-B.1 Models
+FedVS (Li et al., 2023) [^4] Results extracted from graphs. Raw figures were unavailable and the authors did not respond to our inquiries. 81.99 81.83 95.98 98.25
+The feature splits and train-test splits for each dataset are identical to the FedVS paper (Li et al., 2023).
+Note that the datasets here are smaller than MNIST (30000 samples for credit card, 756 for Parkinsons, 1474 for MNIST, 2000 for handwritten).
+Observe that VFL and DVFL both outperform FedVS, which is quantized and constrained only to use polynomial networks (Livni et al., 2014a).
+It is not meaningful to compare these methods under faults, since they are designed for different fault models.
+FedVS is completely tolerant to stragglers, particularly the case where stragglers are bounded in number.
+Like FedVS, this fault model does not affect DVFL, since guests and hosts train independently to each other.
+Conversely, our fault model might cause a catastrophic failure to FedVS: an arbitrary number of participants may crash, which is as if there is a number of stragglers greater than the predetermined system parameter.
+Appendix C Other Failure Modes
+Here, we have considered only crash faults. Other typical failure modes for distributed systems include arbitrary delays, incorrect order, and Byzantine faults.
+Arbitrary Delays and Incorrect Order
+As discussed in the previous section and 3, a line of work exists to partially tackle this problem.
+Typically, methods aimed at handling stragglers either demand a bound on the delay of stragglers or the number thereof.
+In return, methods such as FedVS (Li et al., 2023) can train with no performance loss.
+On the other hand, with bookkeeping like our VFL variants from the main paper, we may observe graceful degradation with stragglers and incorrect order (or no degradation at all in the case of wait, although that also requires a bounded delay).
+Byzantine Faults
+Byzantine faults are the most severe fault model, where messages may be corrupted.
+A possible future direction for DVFL is to incorporate tolerance to Byzantine faults.
+Observe that our SplitNN-buffer strategy is, in a sense, a Byzantine fault, and it is not performant.
+Byzantine faults are not particularly difficult for DVFL during the guest-host phase, since the learning target is not task-relevant.
+But some investigation is required for Byzantine faults during the owner learning phase, which cannot be outright detected as crash faults can.
+Appendix D Models, Hyperparameters and Implementation Details
+D.1 Models
 The “global”, single-participant neural network (see Figure 2) being split is a shallow multi-layer perceptron.
@@
 $W_{h}$ is the output dimension of a single host.
-B.1.1 MNIST
-The “guest” models have 2 layers, which squeeze the flattened $\frac{784}{|\mathcal{G}|}$-long input vector dimension to $\frac{400}{|\mathcal{G}|}$, and then to $\frac{W_{g}}{|\mathcal{G}|}$ with ReLU activation.
-The “host” models squeeze the $W_{g}$-dimensional input to $\frac{W_{g}+3W_{h}}{4}$ and then to $W_{h}$, with Leaky ReLU activation (with negative half-plane slope = 0.01).
-The “owner” model squeezes the $W_{h}\times|\mathcal{H}|$ dimensional input to $W_{h}$ and then to $40$ with a Leaky ReLU activation and finally to a $10$-class prediction with an implicit softmax via the cross entropy loss function.
+D.1.1 MNIST
+Our batch sizes are 64.
+The “guest” models have 2 layers, which squeeze the flattened $\frac{784}{|\mathcal{G}|}$-long input vector dimension to $\frac{400}{|\mathcal{G}|}$ via LeakyReLU, and then to $\frac{W_{g}}{|\mathcal{G}|}$ with ReLU activation.
+The “host” models squeeze the $W_{g}$-dimensional input to $\frac{W_{g}+3W_{h}}{4}$ and then to $W_{h}$ via LeakyReLU, with Leaky ReLU activation (with negative half-plane slope = 0.01).
+The “owner” model squeezes the $W_{h}\times|\mathcal{H}|$ dimensional input to $W_{h}$ and then to $40$ with Leaky ReLU activations and finally to a $10$-class prediction with an implicit softmax via the cross entropy loss function.
 For DVFL experiments, the guest, host, and owner modules are separate objects simulated to reside on separate devices.
@@
 Our experiments in the main paper use $W_{g}=320$ and $W_{h}=160$.
+We run a hyperparameter sweep for each cell.
 $W_{h}$ Test Accuracy (%)
 $W_{g}=200$ $W_{g}=240$ $W_{g}=320$ $W_{g}=400$
-120 96.65 96.93 96.89 96.92
-160 97.07 97.31 97.39 97.15
-200 97.40 97.13 97.11 97.16
+120 97.43 97.28 97.09 97.34
+160 97.60 97.86 97.89 97.69
+200 97.81 97.89 98.15 98.30
 There two competing factors: the number of parameters (a monotonic function of $W_{g}\times W_{h}$) and the compression from host to guest.
@@
 In the main paper, we use $W_{g}=320$ and $W_{h}=160$ for our experiments on MNIST. This choice enables good model performance while also demonstrating host models’ ability to compress the data.
-B.1.2 VFL Datasets
-Similar to MNIST, we use shallow multi-layer perceptrons for our results on vertically partitioned data in Appendix A.
+D.1.2 VFL Datasets
+Similar to MNIST, we use shallow multi-layer perceptrons for our results on vertically partitioned data in the Appendix.
 The model sizes for the guest $g_{i}$, host $h_{k}$ and owner $c$ are summarized in table 6, where $d_{i}$ is the input dimension of guest $g_{i}$.
@@
 Guest Host
-Credit Card $\sum_{i}\lceil\frac{3d_{i}}{4}\rceil$ 10 Leaky ReLU Leaky ReLU
-Parkinsons $\sum_{i}\lfloor{\frac{d_{i}}{4}}\rfloor$ $\frac{\sum_{i}\lfloor{\frac{d_{i}}{4}}\rfloor}{4}$ Leaky ReLU ReLU
+Credit Card $\sum_{i}\left\lceil\frac{3d_{i}}{4}\right\rceil$ 10 Leaky ReLU Leaky ReLU
+Parkinsons $\sum_{i}\left\lfloor\frac{d_{i}}{4}\right\rfloor$ $\frac{\sum_{i}\left\lfloor\frac{d_{i}}{4}\right\rfloor}{4}$ Leaky ReLU ReLU
 CalTech-7 $256\times|\mathcal{G}|$ 256 Leaky ReLU ReLU
-Handwritten $\sum_{i}\lceil\frac{3d_{i}}{4}\rceil$ $3\frac{\sum_{i}\lceil\frac{3d_{i}}{4}\rceil}{4}$ Leaky ReLU ReLU
-B.2 Fault Simulation
+Handwritten $\sum_{i}\left\lceil\frac{3d_{i}}{4}\right\rceil$ $3\frac{\sum_{i}\left\lceil\frac{3d_{i}}{4}\right\rceil}{4}$ Leaky ReLU ReLU
+D.2 Fault Simulation
 We simulate faults based on six hyperparameters:
   1. 1.
-    Communication Faults: During each communication round, for each pair $(g_{j},h_{i})$ we draw a sample from a uniform distribution with support $[0,1]$.
+    Connection Faults: During each connection round, for each pair $(g_{j},h_{i})$ we draw a sample from a uniform distribution with support $[0,1]$.
 We track the current status of each connection using flags.
@@
     In DVFL, we only update $\theta_{j}$ if the guest $g_{j}$ is alive.
-If it is a communication epoch, we only write $g_{j}$’s activations to $\mathbf{B}_{j,i}\forall i\in\{1,2,\dots,|\mathcal{H}|\}$ if the guest is alive.
+If it is a connection epoch, we only write $g_{j}$’s activations to $\mathbf{B}_{j,i}\forall i\in\{1,2,\dots,|\mathcal{H}|\}$ if the guest is alive.
 Similarly, in SplitNN, we only allow a guest to send a forward pass and receive gradients if it is alive.
@@
 When a SplitNN host fails, it does not compute any gradients.
-B.3 Other Hyperparameters
+D.3 Other Hyperparameters
 For all experiments, we perform a hyperparameter search. The values we used for the results presented in this paper are below.
-B.3.1 MNIST
+D.3.1 MNIST
 For guests and DVFL hosts, we use an Adam optimizer with the default $\beta_{1}=0.9,\beta_{2}=0.999$.
-We use a learning rate of $10^{-3}$ and a weight decay of $10^{-5}$.
-For DVFL owners and SplitNN guests we use stochastic gradient descent with a learning rate of $10^{-2}$ and momentum of $0.5$.
+We use a learning rate of $2\times 10^{-3}$ (guest, host) and a weight decay of $10^{-5}$.
+For DVFL owners, we use SGD with a learning rate of $8\times 10^{-2}$.
+For SplitNN guests, we use stochastic gradient descent with a learning rate of $10^{-2}$ and momentum of $0.5$.
 In general, guests were trained for 20 epochs, hosts for 40 and owners for 60.
@@
 In all cases, DVFL owner models and SplitNN host models were trained with early stopping.
-B.3.2 VFL Datasets
-Hyperparameters are presented in Table 6.
-Table 6: Hyperparameters for the results presented in Appendix A.
+D.3.2 VFL Datasets
+Hyperparameters are presented in Table LABEL:morehyper.
+Table 7: Hyperparameters for the results presented in the Appendix.
 Dataset Learning Rate
@@
 Credit Card 1e-3 1e-3 1e-3 1e-3 1e-3
-B.4 Implementation Details
+D.4 Implementation Details
 We simulate parallel training by time multiplexing and using preallocated replay and communication tensors.
-The experiments were run on a variety of hardware, including mobile CPUs (Apple M2) and workstation GPUs (on a single NVIDIA RTX A5000 and single-GPU, and on a single NVIDIA Tesla K40).
-Appendix C Choice of Baseline
-C.1 SplitNN as a Baseline
-In this work we present DVFL, an alternate strategy to VFL.
-We demonstrate its prowess through a DVFL implementation for ANNs.
-Since DVFL is altogether a new strategy to train on vertically partitioned data, it makes sense to benchmark it against the current status quo for VFL with ANNs, which is SplitNN Gupta & Raskar (2018); Ceballos et al.
+The experiments were run on a variety of hardware, including mobile CPUs (Apple Silicon M2) and workstation GPUs (on a single NVIDIA RTX A5000 and on a single NVIDIA Tesla K40).
+Appendix E More Discussion on Related Works
+E.1 Fault-Tolerant Distributed Optimization
+The impact of redundancy for fault tolerance in distributed systems is well studied.
+Liu et al.
+(2021) provide a rigorous analysis in the case of distributed optimization, showing that systems which are $2f$-redundant are approximately resilient to $f$ Byzantine agents.
+The authors also present two fault-tolerant algorithms for horizontally distributed convex optimization.
+Liu (2021); Bouhata et al.
+(2022) extensively review approaches to fault tolerance in distributed learning.
+Most methods require synchronization by an aggregator and incur computational overhead.
+Further, most methods involve manipulating gradients, which forces a choice between convergence rate and tightness.
+For VFL: A limited number of works exist to tackle the related problem of straggling participants.
+Li et al.
+(2023) introduce FedVS, an algorithm for systems where guest models are polynomial networks (Livni et al., 2014a; b).
+FedVS uses a form of secret sharing to introduce redundancy in the input data.
+The redundancy is so arranged such that, when there are at most a fixed number of failed or late guests, timely guests have enough combined information to account for the missing messages.
+In situations where multiple participants might have labels (i.e.
+there is a horizontal federation of label owners), (Xia et al., 2023) exploits redundancy and periodic synchronization to provide straggler resistance.
+One specially designated horizontally federated host is responsible for orchestrating the synchronization between hosts.
+Typically, straggler-resistant algorithms are either designed to tolerate a fixed number of stragglers or a finite duration of lag (i.e.
+a crash would still be catastrophic).
+Further, FedVS and its predecessors (Shi et al., 2022; Chen et al., 2020a; Hu et al., 2019) typically only address faults in the forward pass.
+E.2 Federated Transfer Learning
+Federated Transfer Learning (FTL) is a strategy for systems where there is some overlap in both sample space and feature space.
+The host owns labels and some features (i.e.
+it is an active participant) for entities in its sample space.
+Guests have limited labels for their task, but have some small overlap in feature space and sample space with the host (“aligned” samples).
+Liu et al.
+(2020) present Secure Transfer Federated Learning (STFL), a framework for a single host and a single guest.
+Both parties learn a common latent space using local feature extractors.
+Encodings of the passive participant’s unaligned samples in this feature space is used by the passive participant to learn its task model.
+Feng & Yu (2020); Feng et al.
+(2022) extend this idea to multiple parties.
+Much like DVFL, these methods exploit an unsupervised latent space to learn from data that is not present in the intersection.
+However, unlike DVFL, these methods exploit the existence of common feature space between participants.
+This problem is relevant to robust distributed sensor fusion (Liu, 2021), where there is some redundancy/overlap in the “view” of each participant.
+For the related case of VFL with small $|\mathcal{S}_{\mathrm{guests}}\cap\mathcal{S}_{\mathrm{labels}}|$, Sun et al.
+(2023) introduce few-shot Vertical Federated Learning, where a few epochs of training on the aligned dataset are followed by semi-supervised learning on the unlabelled entities of each client’s dataset.
+To generate labels, guests cluster the gradients they receive from the aligned samples.
+While this enables remarkable model performance, there is unwanted computational overhead required to generate the labels.
+Moreover, this is a form of label inference attack!
+Not only is this a breach of privacy, this clustering could be used downstream by a malicious agent to perform feature inference attacks (see §2.4.3).
+E.3 Greedy and Localized Learning
+Algorithms which update parameters in hidden layers of NNs based on local or auxiliary objectives (as opposed to the global label loss) have recently seen renewed interest.
+In the scope of VFL training, they enable some useful properties for NN training.
+The primary drawbacks of BP in the scope of VFL are the inherent locking synchronization constraints, described in Jaderberg et al.
 (2017).
-SplitNN with a timeout
-In its base form, the SplitNN algorithm does not have any explicit fault handling strategy.
-Implicitly (assuming that the coordination between hosts and guests is interrupt-driven), the hosts use the “wait” strategy.
-Training stops completely if inputs are missing.
-We also consider strategies using “timeout”.
-Timeout allows training to continue, although hampered.
-We handle timeouts by skipping or replacing the values.
-If a message appears after the deadline, it needs to be discarded lest it appear as if it were intended for the next iteration.
-This artificial baseline represents a best-case fault-tolerance for vanilla VFL.
-This way, it gets a fair shot against DVFL.
-Indeed, since DVFL suffers from some information loss (since feedback is very limited), VFL equipped with these strategies may still outperform DVFL at low fault rates.
-Busy-waiting SplitNN
-Under “wait”, training can be made to restart if the failed participant comes back online by busy-waiting viz.
-the repeatedly polling a failed guest.
-If the guests send sample IDs along with their features (and hosts cache messages until they are used), they may figure out which iteration a particular message is intended for.
-This way, no information is lost.
-However, until that failed participant is back online, training is held up.
-This is also a poor strategy from a systems perspective, since it wastes bandwidth and computational resources, especially when there is a message authentication in place.
-If a guest is laggy and the host polls it many times before it repsonds, the cache of messages the host must store will be very large.
-Neither in this case nor the timeout case is there a well-defined strategy to handle backward-pass faults.
-C.2 Other Possible Baselines
-DVFL addresses various VFL issues such as fault tolerance, privacy and sample size with one fell swoop.
-There are various mechanisms built upon standard VFL algorithms (including SplitNN) which aim to address some or the other of these issues.
-FedVS Li et al.
-(2023) is a work that addresses multiple, i.e.
-fault tolerance and privacy.
-FedVS might be a reasonable benchmark for DVFL, and we have presented some results against its paper in Appendix A.
-DVFL outperforms FedVS in perfect conditions.
-In the main paper, we showed that DVFL performance does not degrade much with faults with our generalized fault model.
-On the other hand, FedVS performance is completely fault tolerant, except under a much more stringent fault model.
-Beyond this, FedVS constrains guest models to use a specific type of model– polynomial networks Livni et al.
-(2014a; b).
-Moreover, FedVS only deals with faults during the forward pass.
-FedVS considers a restricted fault model: upto a fixed number of guests can be late to/fail to communicate their activations with the hosts.
-These are also true for most of its predecessors e.g.
-Chen et al.
-(2020a); Hu et al.
-(2019); Shi et al.
-(2022).
-DVFL does not require any such system parameter, and is tolerant to a broader class of faults where any participant can fail at any time, including during backpropagation.
-Appendix D Communication and Space Complexity
-D.1 Periodic Communication
-The owner initiates every round of training as well as communication for all agents.
+In order to update its weights, a layer must wait for the backward signal to reach from the output layer (backward locking).
+In order to begin updating any layer, a full forward pass must first propagate to the end of the network (update locking).
+In order for a layer to evaluate an input to begin with, a forward pass must propagate to it (forward locking).
+These strict locks at the split layer are exactly the failure points for VFL training– forward locking causes training to fail if a guest does not communicate activations, backward locking causes training to fail if the host does not communicate gradients.
+By eschewing the wait to evaluate the global objective, most greedy algorithms are backward unlocked if not update unlocked.
+Belilovsky et al.
+(2020) propose a replay buffer mechanism to partially address forward locking for model parallel training.
+Each agent responsible for training a module uses a replay buffer to store activations from the previous module’s agent.
+During a forward pass, the agent reads an activation from the buffer.
+If the forward signal has propagated through the previous module, it would have written the latest activation to the buffer.
+If not, the previous activation is reused.
+Appendix F Communication and Space Complexity
+F.1 Periodic Communication
+The owner initiates each round of training as well as communication for all agents.
 Since the training of guests and hosts are decoupled, the owner may ask guests to communicate their latent activations (and therefore start host training rounds) at arbitrary intervals.
 In VFL, guest model updates are dependent on host feedback.
-So there is every-shot communication, i.e.
-guests send every single output to the host.
-With one-shot communication such as FedOnce Wu et al.
-(2022), the volume of communication is greatly reduced, at the cost of model performance.
+So there is every-shot communication: guests send every single output to the host.
+With one-shot communication such as FedOnce (Wu et al., 2022), the volume of communication is greatly reduced, at the cost of model performance.
 Equivalently, Belilovsky et al.
@@
 a host wants to train for $k$ iterations and it receives activations of $k$ iterations from the guests.
-This is discussed in more detail in Appendix D.2.
-0055551010101015151515202020202525252596.996.996.996.99797979797.197.197.197.197.297.297.297.297.397.397.397.397.497.497.497.4Number of HostsMNIST Test Accuracy (%)
+This is discussed in more detail in the Appendix F.2.
+005555101010101515151520202020252525259797979797.297.297.297.297.497.497.497.497.697.697.697.697.897.897.897.8Guest training epochs per communication epochMNIST Test Accuracy (%)
 Figure 4: Model performance degrades with an increase in communication period, i.e. more communication is correlated with better performance.
@@
 With 1.536GB, 0.307GB, 0.154GB and 0.077GB outgoing from each guest, the accuracies on the MNIST test set are $97.30\%$, $97.14\%$ $97.08\%$ and $96.84\%$ respectively.
-D.2 Activation Replay Mechanism & its Computational Overhead
-A host reads its model inputs from its activation replay buffer.
-Thanks to this, it can train asynchronously to the guests.
-The host can train independently to the guests as long as the guests have communicated at some prior point of time.
-In fact, that is the case in our experiments– we simulate host models’ training offline, after guest training.
-As discussed in Appendix D.1, this allows control over when the guests communicate.
+F.2 Activation Replay Mechanism & its Computational Overhead
+A host gets its input from its activation replay buffer.
+That allows it to train asynchronously wrt the guests.
+The host can train independently of the guests as long as the guests have communicated earlier.
+Indeed, that is the case in our experiments– we simulate host models’ training offline, after guest training.
+As discussed in Appendix F.1, this allows control over when the guests communicate.
 Another way to look at this is: hosts do not need to train for the same number of iterations as guests communicate messages for.
 This includes the case where guests communicate every iteration, but hosts want to train more (or less) iterations.
-As we detailed in Appendix B.3, we actually implement this for our experiments.
+As we detailed in Appendix D.3, we actually implement this for our experiments.
 While we use 40 epochs to train the entire SplitNN model, we only use 20 epochs to train the DVFL guests.
@@
 […]
+F.3 Empirical Time Complexity
+We compare the following regimes:
+  1. 1.
+    DVFL
+    1. (a)
+      MSE loss
+      • Same setup as main experiments
+      • Guests trained for 20 epochs
+      • Hosts trained for 40 epochs
+      • Owner trained for 60 epochs
+    2. (b)
+      SimCLR loss
+      • Guests trained for 5 epochs
+      • Owner trained for 60 epochs
+      • Guests optimized using Adam with $4\times 10^{-3}$ learning rate, $\beta_{1}=0.9$, $\beta_{2}=0.999$, weight decay $10^{-5}$
+      • $\text{linear}(160,50)\to\text{batchnorm}\to\text{ReLU}\to\text{linear}(50,50)\to\text{batchnorm}\to\text{ReLU}\to\text{linear}(50,50)$ projector
+      • No host
+      • Owner model is equivalent to the host model and owner model of our main experiments in sequence
+      • Owner is optimized using SGD with learning rate $10^{-1}$, momentum 0.9
+      • Batch size of 600.
+  2. 2.
+    VFL
+    1. (a)
+      No localized loss
+      • Same setup as main experiments
+      • System trained for 60 epochs
+We measure the CPU process time on an Apple Silicon M2 CPU and compare the following process + system times:
+  • Forward: time per epoch spent in the forward pass (including loss calculation, if any)
+  • Backward: time per epoch spent in the backward pass (including optimizer step)
+  • Wait: time per epoch spent waiting for gradients
+  • Process: total time the process must be alive
+We do not include communication delay or spent managing tensors, as those are heavily implementation-dependent.
+We also do not count data augmentation time for SimCLR, since that may be precomputed before guest-host training.
+Table 8: DVFL allows guests to train without wait and terminate their processes much sooner than VFL.
+Note that SimCLR requires a large batch size (here: 600), and therefore gets swapped out of memory frequently during training.
+The time reported includes that.
+Strategy
+DVFL (MSE) DVFL (SimCLR) SplitNN
+Quantity Guest forward (s/ep) 2.144 4.872 1.830
+Guest waiting (s/ep) 0.000 0 2.180
+Guest backward (s/ep) 0.650 10.488 0.001
+Host forward (s/ep) 1.748 4.657 0.982
+Host backward (s/ep) 0.806 20.778 0.274
+Owner forward (s/ep) 1.271 1.149
+Owner backward (s/ep) 0.564 2.430
+Guest process (s) 55.883 305.596 80.220
+The accuracies observed are:
+  • SplitNN 97.85%
+  • DVFL (MSE) 97.80%
+  • DVFL (SimCLR) 97.78%
+Accuracies on SimCLR under faults (0 hosts):
+  • Guest
+    1. 1.
+      Rejoin Rate = 1.0: 97.63%
+    2. 2.
+      Rejoin Rate = 0.5: 97.64%
+    3. 3.
+      Rejoin Rate = 0.1: 97.41%
+  • Communication
+    1. 1.
+      Rejoin Rate = 1.0: 97.71%
+    2. 2.
+      Rejoin Rate = 0.5: 97.63%
+    3. 3.
+      Rejoin Rate = 0.1: 97.52%
+The batch size is larger, so losing a single forward pass loses more information.